### v2

# Query complexities of quantum channel discrimination and estimation: A unified approach

Abstract
The goal of quantum channel discrimination and estimation is to determine the identity of an unknown channel from a discrete or continuous set, respectively. The query complexity of these tasks is equal to the minimum number of times one must call an unknown channel to identify it within a desired threshold on the error probability. In this paper, we establish lower bounds on the query complexities of channel discrimination and estimation, in both the parallel and adaptive access models. We do so by establishing new or applying known upper bounds on the squared Bures distance and symmetric logarithmic derivative Fisher information of channels. Phrasing our statements and proofs in terms of isometric extensions of quantum channels allows us to give conceptually simple proofs for both novel and known bounds. We also provide alternative proofs for several established results in an effort to present a consistent and unified framework for quantum channel discrimination and estimation, which we believe will be helpful in addressing future questions in the field.
1 Introduction
1.1 Background
One of the main aims of quantum information theory [Wil17, Hay17, Wat18, Hol19, KW24] is to understand the distinction between the classical and quantum theories of information. Non-classical effects like superposition and entanglement lead to possibilities exceeding what can be achieved in classical information theory alone, with protocols like teleportation [BBC+93], super-dense coding [BW92], and quantum key distribution [BB84, Eke91] being key examples.
Beyond these celebrated protocols, one can also find non-classical features in the tasks of quantum channel discrimination [Sac05, CDP08] and estimation [Fuj01, FI08]. These tasks involve an unknown channel being selected from a predetermined set, and an agent can query the unknown channel multiple times in order to guess or estimate its identity. There are at least two access models for querying the unknown channel, which include the parallel and adaptive access models, as depicted in Figure 1. In channel discrimination, the unknown channel is selected from a finite set, whereas in channel estimation, the unknown channel is selected from a continuous set. As a particularly striking distinction between classical and quantum capabilities in these settings, even if two unitary channels are not perfectly distinguishable with a single query, they always become perfectly distinguishable with a finite number of queries [Ací01]. Additionally, if a unitary channel is selected from a smooth family, one can estimate it with a precision that goes beyond what one can achieve in the classical setting, dubbed the Heisenberg limit [GLM04].
The plethora of rich quantum phenomena that can occur in channel discrimination and estimation has motivated many researchers to pursue and understand this setting for over two decades, so that channel discrimination [Ací01, DFY07, CDP08, DFY09, PW09, Hay09, MPW10, HHLW10, CMW16, WW19, WBHK20, KW21a, SHW22] and channel estimation [Fuj01, SM06, JWD+08, FI08, EdMFD11, Hay11, DDKG12, KDD13, DDM14, YF17, KW21b, ZJ21, KGADD23] now constitute their own subfields within quantum information science research. Additionally, it is understood that there are strong links between the two domains [YF17, KW21b, MKSF+25], given that, roughly speaking, channel estimation can be viewed as discriminating nearby channels.
Figure 1: Depiction of (a) parallel and (b) adaptive strategies for quantum channel discrimination and estimation. In both cases, the unknown channel is queried $n$ times, where $n=3$ in the figure. In discrimination, $\theta\in\{1,2\}$ is from a discrete set, and the goal is for the guess $\hat{\theta}$ to equal $\theta$. In estimation, $\theta\in\Theta\subseteq\mathbb{R}$ is from a continuous set, and the goal is for the guess $\hat{\theta}$ to approximate $\theta$ within some tolerance. In each protocol, $\mathcal{Q}$ is a measurement performed in order to determine the value of $\theta$. In the case of channel discrimination, $\mathcal{Q}$ is a measurement with two outcomes, while in the case of channel estimation, $\mathcal{Q}$ is a measurement with outcomes in $\Theta$.
1.2 Summary of results
In this paper, our main original contributions consist of lower bounds on the error probabilities and query complexities of channel discrimination and estimation. These lower bounds establish fundamental limitations for these tasks, i.e., impossibility statements, and thus, they can be understood as physical limitations placed on any protocol for channel discrimination or estimation, akin to the second law of thermodynamics, the uncertainty principle, or channel capacity theorems. More precisely, the query complexities of quantum channel discrimination and estimation are equal to the number of queries of an unknown channel that are required to determine or estimate its identity, respectively, up to a desired error probability. One can think of query complexity, roughly, as the amount of time one has to wait in order to obtain a desired error probability, and as such, this notion links information theory and complexity theory in a nontrivial way, with query complexity often being used to demonstrate optimality of various algorithms [KKLGT19, HL22].
Although our results here build on many prior results in the domains of channel discrimination and estimation [FI08, DDKG12, KDD13, DDM14, YF17, ZJ21, KGADD23], the lower bounds on query complexities are original to our paper. We should note that the notion of query complexity of channel estimation was introduced recently in [MKSF+25, Section XI] (see also [HKOT23] for work on query complexity of unitary channel estimation); however, the main emphasis of that paper was on establishing a framework for quantum estimation theory in the finite-sample regime, while leaving basic questions regarding channel estimation largely open. Thus, to the best of our knowledge, our lower bounds on the query complexity of channel estimation are the first of their kind in the literature. Furthermore, a number of papers have appeared by now on query complexity of quantum channel discrimination [CDR13, KKLGT19, IM21, HL22, RYCS22, LJW+25, NW25], but none of them have adopted the particular approach taken here in establishing lower bounds.
On the way to establishing these results, we often revisit well known findings in the literature on channel discrimination and estimation. The general approach that we adopt here is to be detailed and comprehensive, even if this amounts to revisiting existing statements in the literature. We believe that this approach has intrinsic value, leading to a consistent and transparent framework and presentation style that should be widely accessible to an audience familiar with the foundations of quantum information theory. We should also note that the technical approach we adopt here is different from the main approach usually adopted in papers on channel discrimination and estimation: indeed, we conduct all of our mathematical developments for these tasks in terms of isometric extensions of channels. In comparison, all prior papers closely related to ours use the Kraus representation of quantum channels for their mathematical developments [FI08, DDKG12, KDD13, DDM14, YF17, ZJ21, KGADD23]. We think an advantage of this approach is that various proofs become easier to follow and have common points, relying on basic properties of isometries and the spectral norm. Additionally, it seems more natural to work with an isometric extension as the key object representing a channel, rather than with Kraus operators, similar to how one manipulates vectors and matrices in linear algebra, rather than manipulating the individual components of the vectors and matrices.
To summarize our results succinctly, we first either recall or establish upper bounds on the squared Bures distance in both the parallel and adaptive settings of channel discrimination (see Theorems 17 and 19), and we do the same for the symmetric logarithmic derivative (SLD) Fisher information for channel estimation (see Theorems 22 and 23). After doing so, we then establish lower bounds on the error probabilities of channel discrimination and estimation in terms of the squared Bures distance and SLD Fisher information (see Corollaries 25, 31, and 34), following a line of thinking adopted in [CDL+25]. By applying the definitions of query complexities for channel estimation and discrimination, we then convert these lower bounds on error probabilities to lower bounds on the query complexities for these tasks (see Corollaries 26, 32, and 34). Finally, we provide various numerical methods for efficiently computing lower bounds on the query complexities of channel discrimination and estimation, in both the parallel and adaptive access models.
Similar to the developments of [YF17], our paper also presents a unified approach to establishing lower bounds on the query complexities of channel discrimination and estimation, making use of a key fact hinted at in [MKSF+25, Section XI] and made more concrete in Section 3.3: lower bounds on the query complexities of channel discrimination imply lower bounds on the query complexities of channel estimation. Indeed, channel estimation can be viewed as attempting to discriminate two nearby channels, and so one would expect the ability to estimate should be limited by the ability to discriminate.
Note on independent work—While finalizing our manuscript, we noticed that the concurrent and independent paper [SDD26] appeared. Similar to the theme of our paper and that of [YF17], Ref. [SDD26] presents a unified approach to quantum channel discrimination and estimation. More specifically, the bound presented in Eqs. (18)–(20) therein are similar in spirit to our bound from Corollary 25.
1.3 Paper organization
The rest of our paper is organized as follows:
  • Section 2 reviews basic concepts and establishes some notation used in the rest of the paper.
  • Section 3 reviews the parallel and adaptive settings of channel discrimination and estimation, defining error probabilities and query complexities in these settings.
  • Section 4 reviews basic aspects of fidelity and Bures distance of states and channels. Although the fidelity and Bures distance were defined some time ago now [Bur69, Uhl76] and have been studied abundantly in the literature, the particular way in which we present these quantities is useful for proving the bounds on channel discrimination and estimation, as presented in later sections.
  • Section 5 reviews the SLD Fisher information of states and channels. In particular, Theorems 12 and 13 by and large date back to [FI08, Theorems 1 and 4]. However, some of the alternative expressions for SLD Fisher information of states, as presented in Theorem 12, appear to be novel, and our proof of Theorem 12 is distinct from the proof of [FI08, Theorem 1]. A similar statement applies to Theorem 13.
  • Section 6 presents upper bounds on the squared Bures distance of channels in both the parallel and adaptive settings of channel discrimination. The upper bound for the parallel setting (Theorem 17) was already established in [YF17, Eq. (23)], but the upper bound for the adaptive setting (Theorem 19) appears to be novel. Regardless, our proofs for these theorems are distinct from that given for [YF17, Eq. (23)] and are consistent with the aforementioned theme: they are established in terms of isometric extensions of quantum channels and rely on basic properties of isometries and the spectral norm.
  • Section 7 reviews upper bounds on the SLD Fisher information of channels in the parallel and adaptive settings of channel estimation. The upper bound in the parallel setting (Theorem 22) was already established in [FI08, Section 4.1], and the upper bound in the adaptive setting (Theorem 23) was established somewhat recently in [KGADD23, Eq. (11)], improving upon an earlier bound from [DDM14, Eq. (9)]. Interestingly, our proofs for Theorems 22 and 23 mirror those given for Theorems 17 and 19, respectively, making use of basic properties of isometries and spectral norms, again consistent with our overall theme. Furthermore, our proofs for Theorems 17 and 19 can be understood as replacing derivatives with finite differences, when compared to the proofs for Theorems 22 and 23, consistent with our other theme of having a unified approach to channel discrimination and estimation.
  • Section 8 presents our main results: lower bounds on the error probabilities and query complexities of channel discrimination and estimation. Ultimately, with the framework in place at this point, the results of this section follow somewhat directly from definitions, reasoning similar to that in [CDL+25, NW25], and the bounds from Theorems 17, 19, 22, and 23.
  • Section 9 is the final technical section of our paper, presenting methods for optimizing bounds on the error probabilities of channel discrimination and estimation by means of semi-definite optimization, or in some cases, semi-definite optimization combined with a grid search. These methods are appealing because they lead to globally optimal numerical solutions of the bounds.
  • We finally conclude in Section 10 by providing a summary of our results and suggestions for future research.
2 Preliminaries
In this section, we establish notation and concepts used throughout the rest of our paper. To begin with, we use the following notation to denote various sets of interest that arise throughout:
Symbol Meaning Definition
$\mathbb{L}$ Set of linear operators
$\mathbb{U}$ Set of unitary operators $\mathbb{U}\coloneqq\left\{U\in\mathbb{L}:U^{{\dagger}}U=UU^{{\dagger}}=I\right\}$
$\mathbb{H}$ Set of Hermitian operators $\mathbb{H}\coloneqq\left\{A\in\mathbb{L}:A=A^{{\dagger}}\right\}$
$\mathbb{D}$ Set of density operators $\mathbb{D}\coloneqq\left\{\rho\in\mathbb{H}:\rho\geq 0,\ \operatorname{Tr}[\rho]=1\right\}$
$\mathbb{P}$ Set of state vectors $\mathbb{P}\coloneqq\left\{|\psi\rangle:\left\||\psi\rangle\right\|=1\right\}$
$\mathbb{B}$ Set of contractions $\mathbb{B}\coloneqq\left\{W\in\mathbb{L}:\left\|W\right\|\leq 1\right\}$
In using this notation, we leave the underlying Hilbert space and its dimension implicit, with the idea being that it should be clear from the context. We also denote an operator $P$ as being positive semi-definite by means of the notation $P\geq 0$, and we denote it as being positive definite by means of the notation $P>0$. In defining the set $\mathbb{P}$ of (pure) state vectors, let us note that $\left\|\cdot\right\|$ denotes the standard Euclidean norm for vectors, and we have left the underlying Hilbert space implicit; however, throughout the paper, we always have $|\psi\rangle\in\mathbb{C}^{d}$ for some $d\in\mathbb{N}$.
An isometry $V$ is an operator satisfying $V^{{\dagger}}V=I$. For an operator $A$, we denote the spectral norm by $\left\|A\right\|\coloneqq\sup_{|\psi\rangle\in\mathbb{P}}\left\|A|\psi\rangle\right\|$. It obeys the following properties for operators $A$ and $B$ and an isometry $V$:
$\left\|A\right\|$ $=\left\|A^{{\dagger}}\right\|=\left\|A^{{\dagger}}A\right\|^{\frac{1}{2}},$ (2.1)
$\left\|AB\right\|$ $\leq\left\|A\right\|\left\|B\right\|,$ (2.2)
$\left\|A\otimes B\right\|$ $=\left\|A\right\|\left\|B\right\|,$ (2.3)
$\left\|A+B\right\|$ $\leq\left\|A\right\|+\left\|B\right\|,$ (2.4)
$\left\|V\right\|$ $=1.$ (2.5)
We also make extensive use of the Schatten norms, defined for $p\geq 1$ as
$\left\|A\right\|_{p}\coloneqq\left(\operatorname{Tr}\!\left[\left|A\right|^{p}\right]\right)^{\frac{1}{p}},$ (2.6)
where $\left|A\right|\coloneqq\sqrt{A^{{\dagger}}A}$. Special cases include the trace norm $\left\|A\right\|_{1}=\operatorname{Tr}\!\left[\left|A\right|\right]$, the Hilbert–Schmidt norm $\left\|A\right\|_{2}=\sqrt{\operatorname{Tr}\!\left[A^{{\dagger}}A\right]}$, and the spectral norm $\left\|A\right\|=\left\|A\right\|_{\infty}$.
When dealing with the SLD Fisher information, we take partial derivatives, which we denote with the shorthand $\partial_{\theta}\equiv\frac{\partial}{\partial\theta}$.
Finally, we make extensive use of the maximally entangled vector
$|\Gamma\rangle\coloneqq\sum_{i}|i\rangle\otimes|i\rangle,$ (2.7)
where $\left\{|i\rangle\right\}_{i}$ is an orthonormal basis, and its associated identities:
$\left(A\otimes I\right)|\Gamma\rangle$ $=\left(I\otimes A^{T}\right)|\Gamma\rangle,$ (2.8)
$\langle\Gamma|\left(A\otimes I\right)|\Gamma\rangle$ $=\operatorname{Tr}\!\left[A\right],$ (2.9)
$\left\|A\right\|_{2}$ $=\left\|\left(A\otimes I\right)|\Gamma\rangle\right\|,$ (2.10)
which hold for a square operator $A$.
3 Quantum channel discrimination and estimation
3.1 Quantum channel discrimination
In the setting of symmetric binary quantum channel discrimination, a channel $\mathcal{N}_{1}$ is selected with probability $p\in\left(0,1\right)$ and a channel $\mathcal{N}_{2}$ is selected with probability $q\equiv 1-p$. The discriminator is allowed to query the unknown channel $n\in\mathbb{N}$ times in an attempt to decide which channel was selected. There are at least two ways for the discriminator to access the channel, called the parallel and the adaptive access models, as depicted in Figure 1.
3.1.1 Parallel and adaptive settings of channel discrimination
In the parallel setting of channel discrimination, the discriminator prepares a state $\rho_{RA_{1}\cdots A_{n}}$, sends the systems $A_{1}\cdots A_{n}$ into $\mathcal{N}_{i}^{\otimes n}$, and performs a measurement on the reference system $R$ and the channel output systems $B_{1}\cdots B_{n}$. The minimum error probability is then given by [KW24, Theorem 5.9]
$p_{e,\|}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n)\coloneqq\frac{1}{2}\left(1-\left\|p\mathcal{N}_{1}^{\otimes n}-q\mathcal{N}_{2}^{\otimes n}\right\|_{\diamond}\right),$ (3.1)
where the diamond norm of a Hermiticity preserving map $\mathcal{M}$ can be defined (for our purposes here) as follows:
$\left\|\mathcal{M}\right\|_{\diamond}\coloneqq\sup_{\psi_{RA}\in\mathbb{P}}\left\|\left(\operatorname{id}_{R}\otimes\mathcal{M}\right)\left(\psi_{RA}\right)\right\|_{1},$ (3.2)
with the reference system $R$ isomorphic to the channel input system $A$.
In the adaptive setting of channel discrimination, the discriminator prepares an initial state $\rho_{RA}$, feeds system $A$ into the first query of the unknown channel $\mathcal{N}_{i}$, then acts on the reference system $R$ and the channel output system with a channel $\mathcal{A}_{1}$, and iterates this process. Overall, a general adaptive $n$-query protocol consists of a tuple $\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}$ of channels. Applying the protocol to $n$ queries of the channel $\mathcal{N}_{i}$ leads to the following channel $\mathcal{P}_{i}^{(n)}$ for $i\in\left\{1,2\right\}$:
$\mathcal{P}_{i}^{(n)}\coloneqq\left(\operatorname{id}\otimes\mathcal{N}_{i}\right)\circ\mathcal{A}_{n-1}\circ\cdots\circ\mathcal{A}_{2}\circ\left(\operatorname{id}\otimes\mathcal{N}_{i}\right)\circ\mathcal{A}_{1}\circ\left(\operatorname{id}\otimes\mathcal{N}_{i}\right).$ (3.3)
The minimum error probability in this case is given by
$p_{e,\operatorname{ad}}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n)\coloneqq\inf_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}\frac{1}{2}\left(1-\left\|p\mathcal{P}_{1}^{(n)}-q\mathcal{P}_{2}^{(n)}\right\|_{\diamond}\right).$ (3.4)
Adopting the shorthand
$p_{e,\|}$ $\equiv p_{e,\|}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n),$ (3.5)
$p_{e,\operatorname{ad}}$ $\equiv p_{e,\operatorname{ad}}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n),$ (3.6)
let us note that the following inequalities hold:
$p_{e,\operatorname{ad}}\leq p_{e,\|}\leq\min\!\left\{p,q\right\}\leq\frac{1}{2}.$ (3.7)
The first inequality follows because a parallel channel discrimination strategy is a special case of an adaptive channel discrimination strategy. The second inequality is a consequence of [ACMT+07, Theorem 1], as remarked in [CDL+25, Eqs. (D1)–(D3)], and the fact that $\min\!\left\{p,q\right\}=\min_{s\in\left[0,1\right]}p^{s}q^{1-s}$. Alternatively, one can understand this inequality as resulting from discarding the channel outputs and simply guessing the channel that has the higher prior probability $p$ or $q$. Discarding the channel in either (3.1) or (3.4) leads to the following expression for the minimum error probability $\frac{1}{2}\left(1-\left|p-q\right|\right)$, which is equal to $\min\{p,q\}$. As such, the following inequalities hold for $p\in\left(0,1\right)$:
$0\leq\frac{p_{e,\operatorname{ad}}\left(1-p_{e,\operatorname{ad}}\right)}{pq}\leq\frac{p_{e,\|}\left(1-p_{e,\|}\right)}{pq}\leq 1,$ (3.8)
because the function $x\mapsto x\left(1-x\right)$ is monotone increasing on the interval $\left[0,\frac{1}{2}\right]$. We can also understand the quantities in (3.8) as being alternative measures of the performance of a channel discrimination protocol, being equal to zero if the channels are perfectly distinguishable and equal to their maximum value of one if the channels are not distinguishable at all, so that the best strategy in this latter case is simply to guess the channel with the higher prior probability $p$ or $q$.
3.1.2 Query complexities of channel discrimination
We can ask about the minimum number of queries needed to achieve an error probability not exceeding a desired threshold. The resulting quantity is known as query complexity and is of chief interest in our paper.
Definition 1.
For $\varepsilon\in\left[0,1\right]$, we define two different notions of query complexity, based on the parallel and adaptive settings of channel discrimination:
$n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\|}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n)\leq\varepsilon\right\},$ (3.9)
$n_{\operatorname{ad}}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\operatorname{ad}}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n)\leq\varepsilon\right\}.$ (3.10)
The quantities above are understood to take the value $\infty$ if the set on the right-hand side is empty.
Intuitively, query complexities give a sense of how long one has to wait in order to achieve a desired error probability in the task of channel discrimination. However, this characterization is somewhat rough, because the definition of query complexity assumes the ability to prepare arbitrary states and perform arbitrary measurements in the parallel setting and it assumes the same ability, in addition to the ability to perform arbitrary channels in the tuple $\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}$, in the adaptive setting, thus ignoring the computational complexity of implementing these operations.
Proposition 7 of [NW25] identifies conditions under which the query complexity of channel discrimination is trivial, i.e., either equal to one, in which case a desired threshold constraint on the error probability can be met with just a single query, or $\infty$, in which case it is impossible to distinguish the channels by using a finite number of queries. In this context, see also [CDL+25, Remark 2]. Here we recall conditions under which the query complexity is equal to one, while noting from (3.7) that two of the conditions from [NW25] and [CDL+25, Remark 2] can actually be merged into a single, simpler condition:
Remark 2 (Trivial cases).
If $\varepsilon\geq\min\!\left\{p,q\right\}$ or if there exists a pure state $\psi_{RA}$ such that
$\operatorname{supp}\!\left(\left(\operatorname{id}\otimes\mathcal{N}_{1}\right)\left(\psi_{RA}\right)\right)\cap\operatorname{supp}\!\left(\left(\operatorname{id}\otimes\mathcal{N}_{2}\right)\left(\psi_{RA}\right)\right)=\emptyset,$ (3.11)
then
$n_{\operatorname{ad}}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)=n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)=1.$ (3.12)
One of the main goals of our paper is to establish lower bounds on the query complexities in (3.9) and (3.10).
3.2 Quantum channel estimation
In the setting of quantum channel estimation, the goal is to estimate a parameter value $\theta\in\Theta$ by querying an unknown channel $\mathcal{N}_{\theta}$ selected from a parameterized family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of channels, where $\Theta\in\mathbb{R}$ is an open set. The estimator is allowed to query the unknown channel $n$ times in order to estimate the value of $\theta$. As with channel discrimination, there are at least two different access models, which include the parallel and adaptive access models, as depicted in Figure 1. In what follows, we detail these settings and recall definitions of the minimax error probabilities and query complexities of channel estimation, following the framework of [MKSF+25, Section XI]. Another main goal of our paper is to establish lower bounds on the query complexities of channel estimation, in both the parallel and adaptive settings.
3.2.1 Parallel setting
Let us begin by considering the parallel setting of channel estimation. An $n$-query parallel estimation protocol begins with the estimator preparing a state $\rho_{RA_{1}\cdots A_{n}}$, feeding the systems $A_{1}\cdots A_{n}$ of this state through the unknown channel $\mathcal{N}_{\theta}^{\otimes n}$, and performing a measurement $\smash{\mathcal{Q}^{\left(n\right)}\coloneqq(Q_{\hat{\theta}}^{\left(n\right)})_{\hat{\theta}\in\Theta}}$ on the reference system $R$ and the channel output systems $B_{1}\cdots B_{n}$. We assume that $\mathcal{Q}^{\left(n\right)}$ forms a positive operator-valued measure (POVM) on $\Theta$ with respect to the Lebesgue measure. The probability of outputting the estimate $\hat{\theta}$, given that $\theta$ is the true parameter value, is as follows:
$\operatorname{Tr}\!\left[Q_{\hat{\theta}}^{\left(n\right)}\left(\operatorname{id}_{R}\otimes\mathcal{N}_{\theta}^{\otimes n}\right)\left(\rho_{RA_{1}\cdots A_{n}}\right)\right],$ (3.13)
and the goal is for $\hat{\theta}$ to be as close to $\theta$ as possible.
Given a tolerance $\delta>0$, the success probability, for identifying an unknown parameter $\theta\in\Theta$ in the family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ by means of a parallel estimation protocol $\mathcal{P}^{\left(n\right)}\equiv\left(\rho_{RA_{1}\cdots A_{n}},\mathcal{Q}^{\left(n\right)}\right)$, is defined as
$p_{s,\|}(\delta,\mathcal{F},\mathcal{P}^{\left(n\right)},\theta)\coloneqq\int d\hat{\theta}\>\operatorname{Tr}\!\left[Q_{\hat{\theta}}^{\left(n\right)}\left(\operatorname{id}_{R}\otimes\mathcal{N}_{\theta}^{\otimes n}\right)\left(\rho_{RA_{1}\cdots A_{n}}\right)\right]w_{\delta}(\hat{\theta}-\theta),$ (3.14)
where $w_{\delta}(\theta^{\prime})$ is a window function, defined as
$w_{\delta}(\theta^{\prime})\coloneqq\begin{cases}1&:\left|\theta^{\prime}\right|\leq\delta\\ 0&:\left|\theta^{\prime}\right|>\delta\end{cases}.$ (3.15)
Intuitively, the success probability $p_{s,\|}(\delta,\mathcal{F},\mathcal{P}^{\left(n\right)},\theta)$ quantifies the probability that the measurement $\mathcal{Q}$ outputs an estimate $\hat{\theta}$ that is $\delta$-close to $\theta$. The error probability is then defined as
$p_{e,\|}(\delta,\mathcal{F},\mathcal{P}^{\left(n\right)},\theta)$ $\coloneqq 1-p_{s,\|}(\delta,\mathcal{F},\mathcal{P}^{\left(n\right)},\theta)$ (3.16)
$=\int_{\hat{\theta}:\left|\hat{\theta}-\theta\right|>\delta}d\hat{\theta}\>\operatorname{Tr}\!\left[Q_{\hat{\theta}}^{\left(n\right)}\left(\operatorname{id}_{R}\otimes\mathcal{N}_{\theta}^{\otimes n}\right)\left(\rho_{RA_{1}\cdots A_{n}}\right)\right].$ (3.17)
The worst-case error probability, for identifying an unknown parameter $\theta\in\Theta$ in the family $\mathcal{F}$ by using a parallel protocol $\mathcal{P}^{\left(n\right)}$, is defined as
$p_{e,\|}(\delta,\mathcal{F},\mathcal{P}^{\left(n\right)})$ $\coloneqq\sup_{\theta\in\Theta}p_{e,\|}(\delta,\mathcal{F},\mathcal{P}^{\left(n\right)},\theta).$ (3.18)
Minimizing the worst-case error probability over every possible parallel protocol $\mathcal{P}^{\left(n\right)}$ leads to the following key figure of merit:
Definition 3 (Minimax error probability for parallel channel estimation).
Given a tolerance $\delta>0$, a parameterized family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of quantum channels, and $n\in\mathbb{N}$, the minimax error probability is defined as
$p_{e,\|}(\delta,\mathcal{F},n)$ $\coloneqq\inf_{\mathcal{P}^{\left(n\right)}}p_{e,\|}(\delta,\mathcal{F},\mathcal{P}^{\left(n\right)})$ (3.19)
$=\inf_{\begin{subarray}{c}\rho_{RA_{1}\cdots A_{n}},\\ \left(Q_{\hat{\theta}}^{\left(n\right)}\right)_{\hat{\theta}\in\Theta}\end{subarray}}\sup_{\theta\in\Theta}\left\{\int_{\hat{\theta}:\left|\hat{\theta}-\theta\right|>\delta}d\hat{\theta}\>\operatorname{Tr}\!\left[Q_{\hat{\theta}}^{\left(n\right)}\left(\operatorname{id}_{R}\otimes\mathcal{N}_{\theta}^{\otimes n}\right)\left(\rho_{RA_{1}\cdots A_{n}}\right)\right]\right\}.$ (3.20)
Definition 4 (Query complexity of parallel channel estimation).
For all $\varepsilon\in\left[0,1\right]$, $\delta>0$, and every parameterized family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of quantum channels, the minimax query complexity of parallel channel estimation is defined as
$n_{\|}^{\star}(\varepsilon,\delta,\mathcal{F})\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\|}(\delta,\mathcal{F},n)\leq\varepsilon\right\}.$ (3.21)
The quantity above is understood to equal $\infty$ if the set on the right-hand side is empty.
As in the case of channel discrimination, the query complexity of channel estimation intuitively captures how long one has to wait, or how many queries one has to make, in order to reach a desired error probability in the task of channel estimation.
3.2.2 Adaptive setting
We now establish similar definitions for the adaptive setting of channel estimation. An $n$-query adaptive channel estimation protocol begins with the estimator preparing an initial state $\rho_{RA}$, feeding system $A$ into the first query of the unknown channel $\mathcal{N}_{\theta}$, then acting on the reference system $R$ and the channel output system with a channel $\mathcal{A}_{1}$, and iterating this process. Overall, a general adaptive $n$-query protocol consists of a tuple $\smash{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}$ of channels. Applying the protocol to $n$ queries of the unknown channel $\mathcal{N}_{\theta}$ leads to the following channel $\mathcal{A}_{\theta}^{(n)}$ for $\theta\in\Theta$:
$\mathcal{A}_{\theta}^{(n)}\coloneqq\left(\operatorname{id}\otimes\mathcal{N}_{\theta}\right)\circ\mathcal{A}_{n-1}\circ\cdots\circ\mathcal{A}_{2}\circ\left(\operatorname{id}\otimes\mathcal{N}_{\theta}\right)\circ\mathcal{A}_{1}\circ\left(\operatorname{id}\otimes\mathcal{N}_{\theta}\right).$ (3.22)
At the end, a measurement $\smash{\mathcal{Q}^{\left(n\right)}\coloneqq(Q_{\hat{\theta}}^{\left(n\right)})_{\hat{\theta}\in\Theta}}$ is performed on the final state. The probability of outputting the estimate $\hat{\theta}$, given that $\theta$ is the true parameter value, is as follows:
$\operatorname{Tr}\!\left[Q_{\hat{\theta}}^{\left(n\right)}\mathcal{A}_{\theta}^{(n)}\!\left(\rho_{RA}\right)\right],$ (3.23)
and the goal is for $\hat{\theta}$ to be as close to $\theta$ as possible.
Given a tolerance $\delta>0$, the error probability, for identifying an unknown parameter $\theta\in\Theta$ in the family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ by means of an adaptive estimation protocol $\mathcal{E}^{\left(n\right)}\equiv\left(\rho_{RA},\left(\mathcal{A}_{i}\right)_{i=1}^{n-1},\mathcal{Q}^{\left(n\right)}\right)$, is defined as
$p_{e,\operatorname{ad}}(\delta,\mathcal{F},\mathcal{E}^{\left(n\right)},\theta)$ $\coloneqq 1-\int d\hat{\theta}\>\operatorname{Tr}\!\left[Q_{\hat{\theta}}^{\left(n\right)}\mathcal{A}_{\theta}^{(n)}\!\left(\rho_{RA}\right)\right]w_{\delta}(\hat{\theta}-\theta)$ (3.24)
$=\int_{\hat{\theta}:\left|\hat{\theta}-\theta\right|>\delta}d\hat{\theta}\>\operatorname{Tr}\!\left[Q_{\hat{\theta}}^{\left(n\right)}\mathcal{A}_{\theta}^{(n)}\!\left(\rho_{RA}\right)\right],$ (3.25)
where $w_{\delta}(\theta^{\prime})$ is the window function defined in (3.15).
The worst-case error probability, for identifying an unknown parameter $\theta\in\Theta$ in the family $\mathcal{F}$ by using an adaptive protocol $\mathcal{E}^{\left(n\right)}$, is defined as
$p_{e,\operatorname{ad}}(\delta,\mathcal{F},\mathcal{E}^{\left(n\right)})$ $\coloneqq\sup_{\theta\in\Theta}p_{e,\operatorname{ad}}(\delta,\mathcal{F},\mathcal{E}^{\left(n\right)},\theta).$ (3.26)
Minimizing the worst-case error probability over every possible adaptive protocol $\mathcal{E}^{\left(n\right)}$ leads to the following key figure of merit:
Definition 5 (Minimax error probability for adaptive channel estimation).
Given a tolerance $\delta>0$, a parameterized family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of quantum channels, and $n\in\mathbb{N}$, the minimax error probability is defined as
$p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)$ $\coloneqq\inf_{\mathcal{E}^{\left(n\right)}}p_{e,\operatorname{ad}}(\delta,\mathcal{F},\mathcal{E}^{\left(n\right)})$ (3.27)
$=\inf_{\begin{subarray}{c}\rho_{RA},\left(\mathcal{A}_{i}\right)_{i=1}^{n-1},\\ \left(Q_{\hat{\theta}}^{\left(n\right)}\right)_{\hat{\theta}\in\Theta}\end{subarray}}\sup_{\theta\in\Theta}\left\{\int_{\hat{\theta}:\left|\hat{\theta}-\theta\right|>\delta}d\hat{\theta}\>\operatorname{Tr}\!\left[Q_{\hat{\theta}}^{\left(n\right)}\mathcal{A}_{\theta}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}.$ (3.28)
Definition 6 (Query complexity of adaptive channel estimation).
For all $\varepsilon\in\left[0,1\right]$, $\delta>0$, and every parameterized family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of quantum channels, the minimax query complexity of adaptive channel estimation is defined as
$n_{\operatorname{ad}}^{\star}(\varepsilon,\delta,\mathcal{F})\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)\leq\varepsilon\right\}.$ (3.29)
The quantity above is understood to equal $\infty$ if the set on the right-hand side is empty.
3.3 Connecting channel estimation and discrimination
Ref. [MKSF+25, Corollary 7] established a non-trivial link between channel estimation and discrimination, particularly relevant for the minimax error formulation discussed in Sections 3.2.1 and 3.2.2. One of the implications of [MKSF+25, Corollary 7], along with the same reasoning that leads to [MKSF+25, Corollary 1], is as follows:
Lemma 7.
For all $\delta>0$ and $n\in\mathbb{N}$ and for every channel family $\mathcal{F}$, the following bounds hold:
$p_{e,\|}(\delta,\mathcal{F},n)$ $\geq\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta:\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}p_{e,\|}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},n\right),$ (3.30)
$p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)$ $\geq\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta:\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}p_{e,\operatorname{ad}}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},n\right).$ (3.31)
Proof.
We provide a simple proof of (3.31) for completeness, which follows Le Cam’s two-point method [Cam73, Yu97]. The proof of (3.30) is analogous. Define the measurement operator $Q_{\left[\theta-\delta,\theta+\delta\right]}^{\left(n\right)}$ as follows:
$Q_{\left[\theta-\delta,\theta+\delta\right]}^{\left(n\right)}\coloneqq\int_{\left|\hat{\theta}-\theta\right|\leq\delta}d\hat{\theta}\>Q_{\hat{\theta}}^{\left(n\right)}.$ (3.32)
Fix $\theta_{1},\theta_{2}\in\Theta$ such that $\left|\theta_{1}-\theta_{2}\right|>2\delta$. Then
$p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)$
$=\inf_{\mathcal{E}^{\left(n\right)}}\sup_{\theta\in\Theta}\left\{\int_{\hat{\theta}:\left|\hat{\theta}-\theta\right|>\delta}d\hat{\theta}\>\operatorname{Tr}\!\left[Q_{\hat{\theta}}^{\left(n\right)}\mathcal{A}_{\theta}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}$ (3.33)
$=\inf_{\mathcal{E}^{\left(n\right)}}\sup_{\theta\in\Theta}\left\{\operatorname{Tr}\!\left[\left(I-Q_{\left[\theta-\delta,\theta+\delta\right]}^{\left(n\right)}\right)\mathcal{A}_{\theta}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}$ (3.34)
$\geq\inf_{\mathcal{E}^{\left(n\right)}}\sup_{\begin{subarray}{c}\theta_{1},\theta_{2}\in\Theta:\\ |\theta_{1}-\theta_{2}|>2\delta\end{subarray}}\left\{\operatorname{Tr}\!\left[\left(I-Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}\right)\mathcal{A}_{\theta_{1}}^{(n)}\!\left(\rho_{RA}\right)\right],\operatorname{Tr}\!\left[\left(I-Q_{\left[\theta_{2}-\delta,\theta_{2}+\delta\right]}^{\left(n\right)}\right)\mathcal{A}_{\theta_{2}}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}$ (3.35)
$\geq\inf_{\mathcal{E}^{\left(n\right)}}\sup_{\begin{subarray}{c}\theta_{1},\theta_{2}\in\Theta:\\ |\theta_{1}-\theta_{2}|>2\delta\end{subarray}}\left\{\operatorname{Tr}\!\left[\left(I-Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}\right)\mathcal{A}_{\theta_{1}}^{(n)}\!\left(\rho_{RA}\right)\right],\operatorname{Tr}\!\left[Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}\mathcal{A}_{\theta_{2}}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}$ (3.36)
$\geq\inf_{\mathcal{E}^{\left(n\right)}}\left\{\frac{1}{2}\operatorname{Tr}\!\left[\left(I-Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}\right)\mathcal{A}_{\theta_{1}}^{(n)}\!\left(\rho_{RA}\right)\right]+\frac{1}{2}\operatorname{Tr}\!\left[Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}\mathcal{A}_{\theta_{2}}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}$ (3.37)
$=p_{e,\operatorname{ad}}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta_{1}},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta_{2}},n\right).$ (3.38)
The second inequality follows because $I-Q_{\left[\theta_{2}-\delta,\theta_{2}+\delta\right]}^{\left(n\right)}\geq Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}$, given that $\left|\theta_{1}-\theta_{2}\right|>2\delta$. Indeed, since $|\theta_{1}-\theta_{2}|>2\delta$, the intervals $\theta_{1}-\delta,\theta_{1}+\delta]$ and $\theta_{2}-\delta,\theta_{2}+\delta]$ are disjoint. Because $\big(Q_{\hat{\theta}}^{(n)}\big)_{\hat{\theta}\in\Theta}$ forms a POVM, it is countably additive on disjoint measurable sets. Hence,
$Q_{[\theta_{1}-\delta,\theta_{1}+\delta]}^{(n)}+Q_{[\theta_{2}-\delta,\theta_{2}+\delta]}^{(n)}=Q_{[\theta_{1}-\delta,\theta_{1}+\delta]\cup[\theta_{2}-\delta,\theta_{2}+\delta]}^{(n)}\leq I,$ (3.39)
where $I$ denotes the identity operator on the final output space. Therefore,
$Q_{[\theta_{1}-\delta,\theta_{1}+\delta]}^{(n)}\leq I-Q_{[\theta_{2}-\delta,\theta_{2}+\delta]}^{(n)},$ (3.40)
which proves the claimed operator inequality. The last equality follows because the optimization ranges over all adaptive discrimination strategies between $\mathcal{N}_{\theta_{1}}$ and $\mathcal{N}_{\theta_{2}}$ with equal prior probabilities, given that every adaptive estimation protocol allows for an arbitrary final POVM on the output system. Hence, the expression coincides with the minimum Bayesian error probability for adaptive discrimination between these two channels. Since the inequality $p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)\geq p_{e,\operatorname{ad}}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta_{1}},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta_{2}},n\right)$ holds for all $\theta_{1},\theta_{2}\in\Theta$ such that $\left|\theta_{1}-\theta_{2}\right|>2\delta$, we conclude (3.31). ∎
Lemma 7 states that the minimax error probabilities for channel estimation, with tolerance $\delta$, are not smaller than the channel discrimination error probabilities for two channels with parameter values that are more than $2\delta$-far from each other. Stated informally, channel estimation is just as difficult as channel discrimination of nearby channels. As such, limitations on channel discrimination provide limitations on channel estimation.
By combining (3.30)–(3.31) with Definitions 1, 3, and 4, we arrive at the following lower bounds on the query complexities of channel estimation in terms of query complexities of channel discrimination, for all $\varepsilon\in\left[0,1\right]$:
$n_{\|}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta:\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}n_{\|}^{\star}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},\varepsilon\right),$ (3.41)
$n_{\operatorname{ad}}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta:\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}n_{\operatorname{ad}}^{\star}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},\varepsilon\right).$ (3.42)
Section 8.2 delves into more detail of our lower bounds on the minimax error probabilities and query complexities of channel estimation.
4 Fidelity and Bures distance of states and channels
4.1 Fidelity and Bures distance of states
In this section, we develop some aspects of fidelity and Bures distance of quantum states. Although these quantities were introduced some time ago [Bur69, Uhl76] and have been extensively studied in the literature, we develop some subtle aspects of these quantities that are typically not considered, yet are necessary in order for us to obtain the tightest query complexity bounds possible in the channel discrimination application, while using available techniques.
To summarize the subtle aspects, the fidelity has a well known characterization in terms of Uhlmann’s theorem [Uhl76], in which it can be written as the squared overlap of two canonical purifications with an optimization over all unitaries acting on the purifying systems of the purifications. However, we can alternatively optimize this same quantity instead over all contractions acting on the purifying system, which is useful in the context of minimax theorems, given that the set of contractions is convex, while the set of unitaries is not. Thus, it is useful to have a form of the fidelity that we can optimize over all contractions. Furthermore, by introducing an extra qubit for the reference system, it is possible to dilate a contraction to a unitary acting on a larger system (called Halmos dilation [Hal50]), which implies that the optimization over contractions acting on the original reference systems is actually equivalent to optimizing over unitaries acting on a larger reference system and vice versa.
After developing these points for the fidelity, we then briefly state how they apply to the Bures distance of states, given the strong connections between Bures distance and fidelity.
4.1.1 Fidelity of states
Recall that the fidelity of states $\rho$ and $\sigma$ is defined as follows [Uhl76]:
$F(\rho,\sigma)\coloneqq\left\|\sqrt{\rho}\sqrt{\sigma}\right\|_{1}^{2}.$ (4.1)
Let us define canonical purifications of $\rho$ and $\sigma$ as follows:
$|\psi^{\rho}\rangle$ $\coloneqq\left(I\otimes\sqrt{\rho}\right)|\Gamma\rangle,$ (4.2)
$|\psi^{\sigma}\rangle$ $\coloneqq\left(I\otimes\sqrt{\sigma}\right)|\Gamma\rangle,$ (4.3)
where $|\Gamma\rangle$ is defined in (2.7).
In Proposition 8 below, we state three different formulas for the root fidelity.
Proposition 8.
For $d$-dimensional states $\rho$ and $\sigma$ with canonical purifications in (4.2) and (4.3), respectively, the following equalities hold:
$\sqrt{F}(\rho,\sigma)$ $=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right]$ (4.4)
$=\sup_{U\in\mathbb{U}}\Re\!\left[\langle\psi^{\rho}|\left(U\otimes I\right)|\psi^{\sigma}\rangle\right]$ (4.5)
$=\sup_{U_{W}\in\mathbb{U}}\Re\!\left[\langle 0|\langle\psi^{\rho}|\left(U_{W}\otimes I\right)|0\rangle|\psi^{\sigma}\rangle\right].$ (4.6)
In (4.6), $|0\rangle|\psi^{\sigma}\rangle,|0\rangle|\psi^{\rho}\rangle\in\mathbb{C}^{2}\otimes\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ and $U_{W}$ acts nontrivially on the first two systems of $\mathbb{C}^{2}\otimes\mathbb{C}^{d}\otimes\mathbb{C}^{d}$.
Proof.
Appealing to the variational form of the trace norm of an operator $A$ as
$\left\|A\right\|_{1}=\sup_{B\in\mathbb{L}}\left\{\left|\operatorname{Tr}\!\left[B^{{\dagger}}A\right]\right|:\left\|B\right\|\leq 1\right\},$ (4.7)
we can write the root fidelity $\sqrt{F}(\rho,\sigma)$ in a variational way as follows:
$\left\|\sqrt{\rho}\sqrt{\sigma}\right\|_{1}$ $=\sup_{W\in\mathbb{B}}\left|\operatorname{Tr}\!\left[\sqrt{\rho}\sqrt{\sigma}W\right]\right|$ (4.8)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\operatorname{Tr}\!\left[\sqrt{\rho}\sqrt{\sigma}W\right]\right]$ (4.9)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\Gamma|\left(I\otimes\sqrt{\rho}\sqrt{\sigma}W\right)|\Gamma\rangle\right]$ (4.10)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\Gamma|\left(W^{T}\otimes\sqrt{\rho}\sqrt{\sigma}\right)|\Gamma\rangle\right]$ (4.11)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\Gamma|\left(I\otimes\sqrt{\rho}\right)\left(W^{T}\otimes I\right)\left(I\otimes\sqrt{\sigma}\right)|\Gamma\rangle\right]$ (4.12)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\psi^{\rho}|\left(W^{T}\otimes I\right)|\psi^{\sigma}\rangle\right]$ (4.13)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right].$ (4.14)
This proves (4.4). The second equality follows because, for $z\in\mathbb{C}$, we can write the polar form $z=re^{i\phi}$, where $r\geq 0$ and $\phi\in\left[0,2\pi\right]$. Then $\left|z\right|=r$ and $\Re\!\left[z\right]=r\cos\phi$, and by tuning a global phase in $W$, we can deduce that the second equality holds. The third equality follows from (2.9) and the fourth from (2.8).
By noting that the trace norm of a square operator $A$ has the following variational characterization as
$\left\|A\right\|_{1}=\sup_{U\in\mathbb{U}}\left|\operatorname{Tr}\!\left[AU\right]\right|,$ (4.15)
we can repeat the analysis above to conclude that
$\left\|\sqrt{\rho}\sqrt{\sigma}\right\|_{1}=\sup_{U\in\mathbb{U}}\Re\left[\langle\psi^{\rho}|\left(U\otimes I\right)|\psi^{\sigma}\rangle\right],$ (4.16)
thus establishing (4.5).
Additionally, we can embed the contraction $W$ into a unitary matrix in a standard way (called Halmos dilation [Hal50]) as follows:
$U_{W}$ $\coloneqq\begin{bmatrix}W&\sqrt{I-WW^{{\dagger}}}\\ \sqrt{I-W^{{\dagger}}W}&-W^{{\dagger}}\end{bmatrix}$ (4.17)
$=|0\rangle\!\langle 0|\otimes W+|0\rangle\!\langle 1|\otimes\sqrt{I-WW^{{\dagger}}}+|1\rangle\!\langle 0|\otimes\sqrt{I-W^{{\dagger}}W}-|1\rangle\!\langle 1|\otimes W^{{\dagger}},$ (4.18)
and observe that
$\left(\langle 0|\otimes I\right)U_{W}\left(|0\rangle\otimes I\right)=W,$ (4.19)
which implies that
$\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle=\langle 0|\langle\psi^{\rho}|\left(U_{W}\otimes I\right)|0\rangle|\psi^{\sigma}\rangle.$ (4.20)
As stated above, $|0\rangle|\psi^{\sigma}\rangle,|0\rangle|\psi^{\rho}\rangle\in\mathbb{C}^{2}\otimes\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ and $U_{W}$ acts nontrivially on the first two systems of $\mathbb{C}^{2}\otimes\mathbb{C}^{d}\otimes\mathbb{C}^{d}$. Using (4.20), we can unify the two formulas for fidelity in (4.14) and (4.16) as follows:
$\left\|\sqrt{\rho}\sqrt{\sigma}\right\|_{1}$ $=\sup_{U_{W}\in\mathbb{U}}\Re\!\left[\langle 0|\langle\psi^{\rho}|\left(U_{W}\otimes I\right)|0\rangle|\psi^{\sigma}\rangle\right]$ (4.21)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right].$ (4.22)
In (4.21), we have used that the optimization can be taken over every unitary $U_{W}$, and not just unitaries of the particular Halmos form in (4.17), because $\left(\langle 0|\otimes I\right)U_{W}\left(|0\rangle\otimes I\right)$ is a contraction for every unitary $U_{W}$. ∎
4.1.2 Bures distance of states
Recall that the Bures distance of states $\rho$ and $\sigma$ is defined as [Bur69]
$d_{B}(\rho,\sigma)$ $\coloneqq\inf_{U\in\mathbb{U}}\left\|\sqrt{\rho}-U\sqrt{\sigma}\right\|_{2}$ (4.23)
$=\inf_{U\in\mathbb{U}}\left\||\psi^{\rho}\rangle-\left(U\otimes I\right)|\psi^{\sigma}\rangle\right\|,$ (4.24)
where $|\psi^{\rho}\rangle$ and $|\psi^{\sigma}\rangle$ are defined in (4.2) and (4.3), respectively, and the second equality follows from (2.10).
Proposition 9.
For $d$-dimensional states $\rho$ and $\sigma$ with canonical purifications in (4.2) and (4.3), respectively, the following equalities hold:
$d_{B}(\rho,\sigma)$ $=\sqrt{2\left(1-\sqrt{F}(\rho,\sigma)\right)}$ (4.25)
$=\inf_{W\in\mathbb{B}}\sqrt{2\left(1-\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right]\right)}$ (4.26)
$=\inf_{U_{W}\in\mathbb{U}}\sqrt{2\left(1-\Re\!\left[\langle 0|\langle\psi^{\rho}|\left(U_{W}\otimes I\right)|0\rangle|\psi^{\sigma}\rangle\right]\right)}$ (4.27)
$=\inf_{U_{W}\in\mathbb{U}}\left\||0\rangle|\psi^{\rho}\rangle-\left(U_{W}\otimes I\right)|0\rangle|\psi^{\sigma}\rangle\right\|.$ (4.28)
In (4.27) and (4.28), $|0\rangle|\psi^{\sigma}\rangle,|0\rangle|\psi^{\rho}\rangle\in\mathbb{C}^{2}\otimes\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ and $U_{W}$ acts nontrivially on the first two systems of $\mathbb{C}^{2}\otimes\mathbb{C}^{d}\otimes\mathbb{C}^{d}$.
Proof.
Consider from (4.24) that
$d_{B}(\rho,\sigma)$ $=\inf_{U\in\mathbb{U}}\left\||\psi^{\rho}\rangle-\left(U\otimes I\right)|\psi^{\sigma}\rangle\right\|$ (4.29)
$=\inf_{U\in\mathbb{U}}\sqrt{\left(\langle\psi^{\rho}|-\langle\psi^{\sigma}|\left(U^{{\dagger}}\otimes I\right)\right)\left(|\psi^{\rho}\rangle-\left(U\otimes I\right)|\psi^{\sigma}\rangle\right)}$ (4.30)
$=\inf_{U\in\mathbb{U}}\sqrt{\langle\psi^{\rho}|\psi^{\rho}\rangle+\langle\psi^{\sigma}|\left(U^{{\dagger}}\otimes I\right)\left(U\otimes I\right)|\psi^{\sigma}\rangle-2\Re\!\left[\langle\psi^{\rho}|\left(U\otimes I\right)|\psi^{\sigma}\rangle\right]}$ (4.31)
$=\inf_{U\in\mathbb{U}}\sqrt{2\left(1-\Re\!\left[\langle\psi^{\rho}|\left(U\otimes I\right)|\psi^{\sigma}\rangle\right]\right)}$ (4.32)
$=\sqrt{2\left(1-\sup_{U\in\mathbb{U}}\Re\!\left[\langle\psi^{\rho}|\left(U\otimes I\right)|\psi^{\sigma}\rangle\right]\right)}$ (4.33)
$=\sqrt{2\left(1-\sqrt{F}(\rho,\sigma)\right)},$ (4.34)
where the last equality follows from Proposition 8. By appealing to (4.21) and (4.22), we can then write the other expressions in (4.25)–(4.28), with (4.28) following by applying the same reasoning as above. ∎
4.2 Bures distance of channels
Now let us consider the case of quantum channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ that have $d_{A}$-dimensional input spaces and $d_{B}$-dimensional output spaces, where $d_{A},d_{B}\in\mathbb{N}$. The Bures distance of channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ is defined as [GLN05]
$d_{B}(\mathcal{N}_{1},\mathcal{N}_{2})\coloneqq\sup_{\rho_{RA}\in\mathbb{D}}d_{B}\!\left(\left(\operatorname{id}\otimes\mathcal{N}_{1}\right)\left(\rho_{RA}\right),\left(\operatorname{id}\otimes\mathcal{N}_{2}\right)\left(\rho_{RA}\right)\right),$ (4.35)
where the supremum is over every bipartite state $\rho_{RA}$ and also implicitly over the dimension of the reference system $R$ (taken to be unbounded in the definition). However, see (4.44) for a simplification of the optimization task.
Canonical Kraus representations for the channels arise from taking spectral decompositions of the Choi operators of these channels [Wil17, Theorem 4.4.1], which are defined for $i\in\left\{1,2\right\}$ as follows:
$\Gamma^{\mathcal{N}_{i}}$ $\coloneqq\left(\operatorname{id}\otimes\mathcal{N}_{i}\right)\left(\Gamma\right),$ (4.36)
$\Gamma$ $\coloneqq|\Gamma\rangle\!\langle\Gamma|.$ (4.37)
Let us denote the canonical Kraus representations by $\left(K_{i}\right)_{i}$ and $\left(L_{j}\right)_{j}$, respectively, and we can take the number $d_{E}$ of them, without loss of generality, to be the same and no more than $d_{A}d_{B}$, so that we simply set $d_{E}=d_{A}d_{B}$. From these, we can construct canonical isometric extensions of the channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ as follows, respectively:
$V_{1}$ $\coloneqq\sum_{i}|i\rangle\otimes K_{i},$ (4.38)
$V_{2}$ $\coloneqq\sum_{j}|j\rangle\otimes L_{j},$ (4.39)
where the orthonormal bases for the environments in each case is the same and the environments have the same dimension. Thus, $V_{i}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ for $i\in\left\{1,2\right\}$, where $d_{E}=d_{A}d_{B}$, and $V_{1}$ and $V_{2}$ are isometries because they satisfy $V_{1}^{{\dagger}}V_{1}=\sum_{i}K_{i}^{{\dagger}}K_{i}=I$ and $V_{2}^{{\dagger}}V_{2}=\sum_{j}L_{j}^{{\dagger}}L_{j}=I$.
The following theorem is strongly related to the developments in [KSW08], in particular Theorem 1 therein.
Theorem 10.
The Bures distance of channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ can be written as follows:
$d_{B}(\mathcal{N}_{1},\mathcal{N}_{2})$ $=\inf_{U_{W}\in\mathbb{U}}\left\|\widetilde{V}_{1}-\left(U_{W}\otimes I_{B}\right)\widetilde{V}_{2}\right\|,$ (4.40)
$=\sqrt{2}\inf_{W\in\mathbb{B}}\left\|I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right\|^{\frac{1}{2}},$ (4.41)
where $\widetilde{V}_{1}$ and $\widetilde{V}_{2}$ are the following isometric extensions of the channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, respectively,
$\widetilde{V}_{1}$ $\coloneqq|0\rangle\otimes V_{1},$ (4.42)
$\widetilde{V}_{2}$ $\coloneqq|0\rangle\otimes V_{2},$ (4.43)
$V_{1}$ and $V_{2}$ are the isometric extensions in (4.38) and (4.39), respectively. Observing that $\widetilde{V}_{i}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{2}\otimes\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$, let us note that the unitary $U_{W}$ acts nontrivially on the first two systems of $\mathbb{C}^{2}\otimes\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$.
Proof.
As a consequence of the data-processing inequality for the Bures distance $d_{B}$ of states, purification, and the Schmidt decomposition theorem, the following equality holds:
$d_{B}(\mathcal{N}_{1},\mathcal{N}_{2})=\sup_{|\psi\rangle_{RA}\in\mathbb{P}}d_{B}\!\left(\left(\operatorname{id}\otimes\mathcal{N}_{1}\right)\left(\psi_{RA}\right),\left(\operatorname{id}\otimes\mathcal{N}_{2}\right)\left(\psi_{RA}\right)\right),$ (4.44)
where $\psi_{RA}\equiv|\psi\rangle\!\langle\psi|_{RA}$ and the reference system $R$ is isomorphic to the channel input system $A$ (see [KW24, Proposition 7.82] for further details). Noting that every pure bipartite state vector $|\psi\rangle_{RA}$ can be written as follows:
$|\psi\rangle_{RA}=\left(U_{R}\otimes\sqrt{\rho}\right)|\Gamma\rangle$ (4.45)
where $U_{R}$ is a unitary and $\rho$ is a density operator, we can take canonical purifications of $\left(\operatorname{id}\otimes\mathcal{N}_{1}\right)\left(\psi_{RA}\right)$ and $\left(\operatorname{id}\otimes\mathcal{N}_{2}\right)\left(\psi_{RA}\right)$ to be as follows, respectively:
$V_{1}|\psi\rangle_{RA}$ $=\left(U_{R}\otimes V_{1}\sqrt{\rho}\right)|\Gamma\rangle=\sum_{i}|i\rangle_{E}\otimes\left(U_{R}\otimes K_{i}\sqrt{\rho}\right)|\Gamma\rangle,$ (4.46)
$V_{2}|\psi\rangle_{RA}$ $=\left(U_{R}\otimes V_{2}\sqrt{\rho}\right)|\Gamma\rangle=\sum_{j}|j\rangle_{E}\otimes\left(U_{R}\otimes L_{j}\sqrt{\rho}\right)|\Gamma\rangle.$ (4.47)
By appealing to (4.25)–(4.28), we can write the Bures distance of channels as follows:
$d_{B}(\mathcal{N}_{1},\mathcal{N}_{2})$ $=\sup_{\left|\psi\right\rangle_{RA}\in\mathbb{P}}\sqrt{2\left(1-\sqrt{F}\!\left(\left(\operatorname{id}\otimes\mathcal{N}_{1}\right)\left(\psi_{RA}\right),\left(\operatorname{id}\otimes\mathcal{N}_{2}\right)\left(\psi_{RA}\right)\right)\right)}$ (4.48)
$=\sup_{\left|\psi\right\rangle_{RA}\in\mathbb{P}}\inf_{W\in\mathbb{B}}\sqrt{2\left(1-\Re\!\left[\langle\psi|_{RA}V_{1}^{{\dagger}}\left(W_{E}\otimes I_{RB}\right)V_{2}|\psi\rangle_{RA}\right]\right)},$ (4.49)
where we have denoted the channel output system by $B$ and the environment system by $E$. Now consider that
$\langle\psi|_{RA}V_{1}^{{\dagger}}\left(W_{E}\otimes I_{RB}\right)V_{2}|\psi\rangle_{RA}$
$=\langle\psi|_{RA}\left(I_{R}\otimes V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right)|\psi\rangle_{RA}$ (4.50)
$=\langle\Gamma|\left(U_{R}^{{\dagger}}\otimes\sqrt{\rho}\right)\left(I_{R}\otimes V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right)\left(U_{R}\otimes\sqrt{\rho}\right)|\Gamma\rangle$ (4.51)
$=\langle\Gamma|\left(U_{R}^{{\dagger}}U_{R}\otimes\sqrt{\rho}\right)V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\sqrt{\rho}|\Gamma\rangle$ (4.52)
$=\langle\Gamma|\left(I_{R}\otimes\sqrt{\rho}\ V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\sqrt{\rho}\right)|\Gamma\rangle$ (4.53)
$=\operatorname{Tr}\!\left[\sqrt{\rho}\ V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\sqrt{\rho}\right]$ (4.54)
$=\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right].$ (4.55)
Then
$\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\inf_{W\in\mathbb{B}}\sqrt{2\left(1-\Re\!\left[\langle\psi|_{RA}V_{1}^{{\dagger}}\left(W_{E}\otimes I_{RB}\right)V_{2}|\psi\rangle_{RA}\right]\right)}$
$=\sup_{\rho\in\mathbb{D},U\in\mathbb{U}}\inf_{W\in\mathbb{B}}\sqrt{2\left(1-\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.56)
$=\sup_{\rho\in\mathbb{D}}\inf_{W\in\mathbb{B}}\sqrt{2\left(1-\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.57)
$=\sqrt{2\left(1-\inf_{\rho\in\mathbb{D}}\sup_{W\in\mathbb{B}}\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.58)
$=\sqrt{2\left(1-\sup_{W\in\mathbb{B}}\inf_{\rho\in\mathbb{D}}\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.59)
$=\inf_{W\in\mathbb{B}}\sup_{\rho\in\mathbb{D}}\sqrt{2\left(1-\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.60)
$=\inf_{W\in\mathbb{B}}\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\sqrt{2\left(1-\Re\!\left[\langle\psi|_{RA}V_{1}^{{\dagger}}\left(W_{E}\otimes I_{RB}\right)V_{2}|\psi\rangle_{RA}\right]\right)}$ (4.61)
$=\inf_{U_{W}\in\mathbb{U}}\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\sqrt{2\left(1-\Re\!\left[\langle 0|\langle\psi|_{RA}V_{1}^{{\dagger}}\left(U_{W}\otimes I_{RB}\right)|0\rangle V_{2}|\psi\rangle_{RA}\right]\right)}$ (4.62)
$=\inf_{U_{W}\in\mathbb{U}}\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\left\||0\rangle V_{1}|\psi\rangle_{RA}-\left(U_{W}\otimes I_{RB}\right)|0\rangle V_{2}|\psi\rangle_{RA}\right\|$ (4.63)
$=\inf_{U_{W}\in\mathbb{U}}\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\left\|\widetilde{V}_{1}|\psi\rangle_{RA}-\left(U_{W}\otimes I_{RB}\right)\widetilde{V}_{2}|\psi\rangle_{RA}\right\|$ (4.64)
$=\inf_{U_{W}\in\mathbb{U}}\left\|I_{R}\otimes\widetilde{V}_{1}-\left(U_{W}\otimes I_{RB}\right)\left(I_{R}\otimes\widetilde{V}_{2}\right)\right\|$ (4.65)
$=\inf_{U_{W}\in\mathbb{U}}\left\|I_{R}\otimes\left(\widetilde{V}_{1}-\left(U_{W}\otimes I_{B}\right)\widetilde{V}_{2}\right)\right\|$ (4.66)
$=\inf_{U_{W}\in\mathbb{U}}\left\|\widetilde{V}_{1}-\left(U_{W}\otimes I_{B}\right)\widetilde{V}_{2}\right\|,$ (4.67)
The equality in (4.59) follows from the minimax theorem: indeed, the set of density operators is convex and compact, the set of contractions is convex and compact, the function
$\rho\mapsto\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]$ (4.68)
is linear in $\rho$ (and thus convex in $\rho$), and the function
$W\mapsto\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]$ (4.69)
is real linear in $W$ (and thus concave in $W$). The equality in (4.61) follows from (4.55). Additionally, the equality in (4.64) follows from defining $\widetilde{V}_{i}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{2}\otimes\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ as the following isometry, for $i\in\left\{1,2\right\}$:
$\widetilde{V}_{i}|\varphi\rangle_{A}\coloneqq|0\rangle\otimes V_{i}|\varphi\rangle_{A}.$ (4.70)
That is, these isometries can be written as in (4.42)–(4.43). The final equality in (4.67) follows because the spectral norm is multiplicative on tensor products (i.e., $\left\|A\otimes B\right\|=\left\|A\right\|\left\|B\right\|$) and $\left\|I\right\|=1$. ∎
Remark 11.
Relevant for our applications in channel discrimination, we can alternatively write the overlap term in Theorem 10 in terms of the Kraus operators $\left(K_{i}\right)_{i}$ and $\left(L_{j}\right)_{j}$ of channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, respectively, as follows:
$V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}$ $=\left(\sum_{i}\langle i|\otimes K_{i}^{{\dagger}}\right)\left(W_{E}\otimes I_{B}\right)\left(\sum_{j}|j\rangle\otimes L_{j}\right)$ (4.71)
$=\sum_{i,j}\langle i|W|j\rangle\otimes K_{i}^{{\dagger}}L_{j}$ (4.72)
$=\sum_{i,j}W_{i,j}K_{i}^{{\dagger}}L_{j},$ (4.73)
where we defined $W_{i,j}\coloneqq\langle i|W|j\rangle$. This allows for connecting to the formulations in [YF17].
5 SLD Fisher information of states and channels
5.1 SLD Fisher information of states
In this section, we provide alternative formulas for the symmetric logarithmic derivative (SLD) Fisher information of a smooth family $\left(\rho_{\theta}\right)_{\theta\in\Theta}$ of positive definite states. These formulations are related to those given in [FI08, Theorem 1], but we expand on the formulation given there and provide an alternative proof. To begin with, let us recall that the SLD Fisher information of a smooth family $\left(\rho_{\theta}\right)_{\theta\in\Theta}$ of positive definite states can be written as follows [SK20, Eq. (88)]:
$I_{F}\!\left(\theta;\left(\rho_{\theta}\right)_{\theta}\right)=\sum_{\ell,m}\frac{2}{\lambda_{\ell}+\lambda_{m}}\operatorname{Tr}\!\left[\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\left(\partial_{\theta}\rho_{\theta}\right)\right],$ (5.1)
where $\partial_{\theta}\equiv\frac{\partial}{\partial\theta}$,
$\rho_{\theta}=\sum_{k}\lambda_{k}\Pi_{k}$ (5.2)
is the spectral decomposition of $\rho_{\theta}$, with the dependence of each eigenvalue $\lambda_{k}$ and each eigenprojection $\Pi_{k}$ on $\theta$ suppressed in the notation. It is well known that the following equalities hold, relating Bures distance to SLD Fisher information [Hü92]:
$d_{B}^{2}\!\left(\rho_{\theta},\rho_{\theta+\delta}\right)$ $=\frac{\delta^{2}}{4}I_{F}\!\left(\theta;\left(\rho_{\theta}\right)_{\theta}\right)+o(\delta^{2})$ (5.3)
$I_{F}\!\left(\theta;\left(\rho_{\theta}\right)_{\theta}\right)$ $=2\left.\frac{\partial^{2}}{\partial\delta^{2}}d_{B}^{2}\!\left(\rho_{\theta},\rho_{\theta+\delta}\right)\right|_{\delta=0}$ (5.4)
$=4\lim_{\delta\to 0}\frac{d_{B}^{2}\!\left(\rho_{\theta},\rho_{\theta+\delta}\right)}{\delta^{2}}.$ (5.5)
Theorem 12.
For a smooth family $\left(\rho_{\theta}\right)_{\theta\in\Theta}$ of positive definite states, the following equality holds:
$\frac{1}{4}I_{F}\!\left(\theta;\left(\rho_{\theta}\right)_{\theta}\right)$ $=\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\partial_{\theta}\left(\left(U_{\theta}\otimes I\right)|\psi_{\theta}\rangle\right)\right\|^{2}$ (5.6)
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)-i\sqrt{\rho_{\theta}}H_{\theta}\right\|_{2}^{2}$ (5.7)
$=\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]-\operatorname{Tr}\!\left[\left(H_{\theta}^{\star}\right)^{2}\rho_{\theta}\right],$ (5.8)
where $|\psi_{\theta}\rangle\coloneqq\left(I\otimes\sqrt{\rho_{\theta}}\right)|\Gamma\rangle$ is the canonical purification of $\rho_{\theta}$ and the optimizations in (5.6) and (5.7) are over every smooth family $\left(U_{\theta}\right)_{\theta\in\Theta}$ of unitary operators and every smooth family $\left(H_{\theta}\right)_{\theta\in\Theta}$ of Hermitian operators, respectively. The optimal choice $H_{\theta}^{\star}$ in (5.7) satisfies
$\rho_{\theta}H_{\theta}^{\star}+H_{\theta}^{\star}\rho_{\theta}=i\left[\partial_{\theta}\sqrt{\rho_{\theta}},\sqrt{\rho_{\theta}}\right]$ (5.9)
and is thus given by
$H_{\theta}^{\star}$ $\coloneqq i\int_{0}^{\infty}ds\ e^{-s\rho_{\theta}}\left[\partial_{\theta}\sqrt{\rho_{\theta}},\sqrt{\rho_{\theta}}\right]e^{-s\rho_{\theta}}.$ (5.10)
$=-i\sum_{\ell,m}\frac{\lambda_{\ell}^{\frac{1}{2}}-\lambda_{m}^{\frac{1}{2}}}{\left(\lambda_{\ell}+\lambda_{m}\right)\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)}\,\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m},$ (5.11)
where we take the spectral decomposition of $\rho_{\theta}$ as in (5.2).
Proof.
Consider that
$\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\partial_{\theta}\left(\left(U_{\theta}\otimes I\right)|\psi_{\theta}\rangle\right)\right\|^{2}$ $=\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\left(\left(\partial_{\theta}U_{\theta}\right)\otimes I\right)|\psi_{\theta}\rangle+\left(U_{\theta}\otimes I\right)|\partial_{\theta}\psi_{\theta}\rangle\right\|^{2}$ (5.12)
$=\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|U_{\theta}^{{\dagger}}\left[\left(\left(\partial_{\theta}U_{\theta}\right)\otimes I\right)|\psi_{\theta}\rangle+\left(U_{\theta}\otimes I\right)|\partial_{\theta}\psi_{\theta}\rangle\right]\right\|^{2}$ (5.13)
$=\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\left(U_{\theta}^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)\otimes I\right)|\psi_{\theta}\rangle+|\partial_{\theta}\psi_{\theta}\rangle\right\|^{2}$ (5.14)
$=\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\left|\partial_{\theta}\psi_{\theta}\right\rangle+\left(U_{\theta}^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)\otimes I\right)|\psi_{\theta}\rangle\right\|^{2}.$ (5.15)
We claim that $\left(H_{\theta}\right)_{\theta\in\Theta}$, where $iU_{\theta}^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)\equiv H_{\theta}\in\mathbb{H}$, is a smooth family of Hermitian operators. Indeed, by setting $U_{\theta}=e^{-iG_{\theta}}$, we conclude from Duhamel’s formula (see, e.g., [Wil25, Proposition 47]) that
$i\left(U_{\theta}\right)^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)$ $=ie^{iG_{\theta}}\left(\partial_{\theta}e^{-iG_{\theta}}\right)$ (5.16)
$=ie^{iG_{\theta}}\int_{0}^{1}dt\ e^{-iG_{\theta}t}\left(\partial_{\theta}\left(-iG_{\theta}\right)\right)e^{-iG_{\theta}\left(1-t\right)}$ (5.17)
$=\int_{0}^{1}dt\ e^{iG_{\theta}\left(1-t\right)}\left(\partial_{\theta}G_{\theta}\right)e^{-iG_{\theta}\left(1-t\right)},$ (5.18)
which is evidently Hermitian. To solve for $U_{\theta}$ from $H_{\theta}$, we can solve the following matrix differential equation for $U_{\theta}$:
$H_{\theta}=i\left(U_{\theta}\right)^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)\qquad\Longleftrightarrow\qquad-iU_{\theta}H_{\theta}=\partial_{\theta}U_{\theta},$ (5.19)
which is equivalent to the time-dependent Schrödinger equation and well known to have a solution given by the following Dyson series [RS75]:
$U_{\theta}=I+\sum_{k=1}^{\infty}\left(-i\right)^{k}\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}\cdots\int_{0}^{t_{k-1}}dt_{k}\,H(t_{1})H(t_{2})\cdots H(t_{k}).$ (5.20)
This Dyson series is often abbreviated as the time-ordered exponential $\mathcal{T}\exp\!\left(-i\int_{0}^{t}d\tau\,H(\tau)\right)$. Continuing, we make the substitution $U_{\theta}^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)\to-iH_{\theta}$ and find that
$\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\left|\partial_{\theta}\psi_{\theta}\right\rangle+\left(U_{\theta}^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)\otimes I\right)|\psi_{\theta}\rangle\right\|^{2}$
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\left|\partial_{\theta}\psi_{\theta}\right\rangle-i\left(H_{\theta}\otimes I\right)|\psi_{\theta}\rangle\right\|^{2}$ (5.21)
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\left(I\otimes\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)\right)|\Gamma\rangle-i\left(H_{\theta}\otimes\sqrt{\rho_{\theta}}\right)|\Gamma\rangle\right\|^{2}$ (5.22)
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\left(I\otimes\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)\right)|\Gamma\rangle-i\left(I\otimes\sqrt{\rho_{\theta}}H_{\theta}^{T}\right)|\Gamma\rangle\right\|^{2}$ (5.23)
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\left(I\otimes\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)-i\left(I\otimes\sqrt{\rho_{\theta}}H_{\theta}\right)\right)|\Gamma\rangle\right\|^{2}$ (5.24)
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)-i\sqrt{\rho_{\theta}}H_{\theta}\right\|_{2}^{2}.$ (5.25)
At this point, observe that the function $H_{\theta}\mapsto\left\|\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)-i\sqrt{\rho_{\theta}}H_{\theta}\right\|_{2}^{2}$ is convex in $H_{\theta}$, because it is a composition of a convex function (the Schatten $2$-norm) and a convex and non-decreasing function (the square function $x\mapsto x^{2}$). Since our goal is to minimize this expression, convexity in $H_{\theta}$ implies that a local minimum is a global minimum. Let us expand the objective function as follows:
$\left\|\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)-i\sqrt{\rho_{\theta}}H_{\theta}\right\|_{2}^{2}$ $=\operatorname{Tr}\!\left[\left(\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)-i\sqrt{\rho_{\theta}}H_{\theta}\right)^{{\dagger}}\left(\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)-i\sqrt{\rho_{\theta}}H_{\theta}\right)\right]$ (5.26)
$=\operatorname{Tr}\!\left[\left(\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)+iH_{\theta}\sqrt{\rho_{\theta}}\right)\left(\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)-i\sqrt{\rho_{\theta}}H_{\theta}\right)\right]$ (5.27)
$=\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]-i\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)\sqrt{\rho_{\theta}}H_{\theta}\right]$
$\qquad+i\operatorname{Tr}\!\left[H_{\theta}\sqrt{\rho_{\theta}}\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)\right]+\operatorname{Tr}\!\left[H_{\theta}^{2}\rho_{\theta}\right]$ (5.28)
$=\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]+\operatorname{Tr}\!\left[H_{\theta}^{2}\rho_{\theta}\right]-i\operatorname{Tr}\!\left[\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right),\sqrt{\rho_{\theta}}\right]H_{\theta}\right].$ (5.29)
We can then take the matrix derivative of (5.29) with respect to $H_{\theta}$ and find that
$\frac{\partial}{\partial H_{\theta}}\left(\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]+\operatorname{Tr}\!\left[H_{\theta}^{2}\rho_{\theta}\right]-i\operatorname{Tr}\!\left[\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right),\sqrt{\rho_{\theta}}\right]H_{\theta}\right]\right)\\ =\rho_{\theta}H_{\theta}+H_{\theta}\rho_{\theta}-i\left[\partial_{\theta}\sqrt{\rho_{\theta}},\sqrt{\rho_{\theta}}\right],$ (5.30)
where we used the facts that $\frac{\partial}{\partial X}\operatorname{Tr}\!\left[AX^{2}\right]=AX+XA$ and $\frac{\partial}{\partial X}\operatorname{Tr}\!\left[AX\right]=A$. Setting the matrix derivative equal to zero and rewriting implies that the optimal $H_{\theta}^{\star}$ satisfies the following equation:
$\rho_{\theta}H_{\theta}^{\star}+H_{\theta}^{\star}\rho_{\theta}=i\left[\partial_{\theta}\sqrt{\rho_{\theta}},\sqrt{\rho_{\theta}}\right].$ (5.31)
By plugging (5.31) into (5.29), observe that the objective function can be written as
$\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]+\operatorname{Tr}\!\left[\left(H_{\theta}^{\star}\right)^{2}\rho_{\theta}\right]-i\operatorname{Tr}\!\left[\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right),\sqrt{\rho_{\theta}}\right]H_{\theta}^{\star}\right]$
$=\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]+\operatorname{Tr}\!\left[\left(H_{\theta}^{\star}\right)^{2}\rho_{\theta}\right]-\operatorname{Tr}\!\left[\left(\rho_{\theta}H_{\theta}^{\star}+H_{\theta}^{\star}\rho_{\theta}\right)H_{\theta}^{\star}\right]$ (5.32)
$=\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]+\operatorname{Tr}\!\left[\left(H_{\theta}^{\star}\right)^{2}\rho_{\theta}\right]-2\operatorname{Tr}\!\left[\left(H_{\theta}^{\star}\right)^{2}\rho_{\theta}\right]$ (5.33)
$=\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]-\operatorname{Tr}\!\left[\left(H_{\theta}^{\star}\right)^{2}\rho_{\theta}\right].$ (5.34)
By employing (5.31), this formula can alternatively be written as
$\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]-\frac{i}{2}\operatorname{Tr}\!\left[\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right),\sqrt{\rho_{\theta}}\right]H_{\theta}^{\star}\right].$ (5.35)
The equation in (5.31) has the form of a Sylvester–Lyapunov equation, which is well known to have the following solution [Lan70, Theorem 5]:
$H_{\theta}^{\star}$ $=i\int_{0}^{\infty}ds\ e^{-s\rho_{\theta}}\left[\partial_{\theta}\sqrt{\rho_{\theta}},\sqrt{\rho_{\theta}}\right]e^{-s\rho_{\theta}}$ (5.36)
$=i\sum_{\ell,m}\frac{1}{\lambda_{\ell}+\lambda_{m}}\Pi_{\ell}\left[\partial_{\theta}\sqrt{\rho_{\theta}},\sqrt{\rho_{\theta}}\right]\Pi_{m},$ (5.37)
where we employed the integral $\int_{0}^{\infty}ds\ e^{-s\left(x+y\right)}=\frac{1}{x+y}$, holding for $x,y>0$, as well as the spectral decomposition of $\rho_{\theta}$. Now recall that (see, e.g., [Wil25, Theorem 42])
$\partial_{\theta}\sqrt{\rho_{\theta}}$ $=\sum_{\ell,m}f_{x^{\frac{1}{2}}}^{[1]}(\lambda_{\ell},\lambda_{m})\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}$ (5.38)
$=\sum_{\ell,m}\frac{1}{\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}}\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m},$ (5.39)
where we used the first divided difference of the square-root function, defined for $x,y>0$ as
$f_{x^{\frac{1}{2}}}^{[1]}(x,y)\coloneqq\begin{cases}\frac{1}{2x^{\frac{1}{2}}}&:x=y\\ \frac{x^{\frac{1}{2}}-y^{\frac{1}{2}}}{x-y}&:x\neq y\end{cases}.$ (5.40)
This further implies that
$\left[\partial_{\theta}\sqrt{\rho_{\theta}},\sqrt{\rho_{\theta}}\right]$ $=\left[\sum_{r,s}\frac{1}{\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}}\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s},\sqrt{\rho_{\theta}}\right]$ (5.41)
$=\sum_{r,s}\frac{1}{\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}}\left[\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s},\sqrt{\rho_{\theta}}\right]$ (5.42)
$=\sum_{r,s}\frac{1}{\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}}\left(\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s}\sqrt{\rho_{\theta}}-\sqrt{\rho_{\theta}}\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s}\right)$ (5.43)
$=\sum_{r,s}\left(\frac{\lambda_{s}^{\frac{1}{2}}-\lambda_{r}^{\frac{1}{2}}}{\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}}\right)\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s},$ (5.44)
and
$H_{\theta}^{\star}$ $=i\sum_{\ell,m}\frac{1}{\lambda_{\ell}+\lambda_{m}}\Pi_{\ell}\left[\partial_{\theta}\sqrt{\rho_{\theta}},\sqrt{\rho_{\theta}}\right]\Pi_{m}$ (5.45)
$=i\sum_{\ell,m}\frac{1}{\lambda_{\ell}+\lambda_{m}}\Pi_{\ell}\sum_{r,s}\left(\frac{\lambda_{s}^{\frac{1}{2}}-\lambda_{r}^{\frac{1}{2}}}{\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}}\right)\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s}\Pi_{m}$ (5.46)
$=i\sum_{\ell,m,r,s}\frac{\lambda_{s}^{\frac{1}{2}}-\lambda_{r}^{\frac{1}{2}}}{\left(\lambda_{\ell}+\lambda_{m}\right)\left(\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}\right)}\,\Pi_{\ell}\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s}\Pi_{m}$ (5.47)
$=i\sum_{\ell,m}\frac{\lambda_{m}^{\frac{1}{2}}-\lambda_{\ell}^{\frac{1}{2}}}{\left(\lambda_{\ell}+\lambda_{m}\right)\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)}\,\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}$ (5.48)
$=-i\sum_{\ell,m}\frac{\left(\lambda_{\ell}^{\frac{1}{2}}-\lambda_{m}^{\frac{1}{2}}\right)}{\left(\lambda_{\ell}+\lambda_{m}\right)\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)}\,\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}.$ (5.49)
Then consider that
$\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]-\frac{i}{2}\operatorname{Tr}\!\left[\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right),\sqrt{\rho_{\theta}}\right]H_{\theta}^{\star}\right]$
$=\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]-\frac{i}{2}\operatorname{Tr}\!\left[\begin{array}[]{c}\left(\sum_{r,s}\left(\frac{\lambda_{s}^{\frac{1}{2}}-\lambda_{r}^{\frac{1}{2}}}{\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}}\right)\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s}\right)\times\\ \left(-i\sum_{\ell,m}\frac{\left(\lambda_{\ell}^{\frac{1}{2}}-\lambda_{m}^{\frac{1}{2}}\right)}{\left(\lambda_{\ell}+\lambda_{m}\right)\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)}\,\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\right)\end{array}\right]$ (5.52)
$=\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]-\frac{1}{2}\operatorname{Tr}\!\left[\begin{array}[]{c}\left(\sum_{r,s}\left(\frac{\lambda_{s}^{\frac{1}{2}}-\lambda_{r}^{\frac{1}{2}}}{\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}}\right)\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s}\right)\times\\ \left(\sum_{\ell,m}\frac{\left(\lambda_{\ell}^{\frac{1}{2}}-\lambda_{m}^{\frac{1}{2}}\right)}{\left(\lambda_{\ell}+\lambda_{m}\right)\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)}\,\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\right)\end{array}\right].$ (5.55)
Furthermore,
$\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]$ $=\operatorname{Tr}\!\left[\left(\sum_{\ell,m}\frac{1}{\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}}\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\right)^{2}\right]$ (5.56)
$=\sum_{\ell,m,r,s}\left(\frac{1}{\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}}\right)\left(\frac{1}{\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}}\right)\operatorname{Tr}\!\left[\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s}\right]$ (5.57)
$=\sum_{\ell,m}\frac{1}{\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)^{2}}\operatorname{Tr}\!\left[\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\left(\partial_{\theta}\rho_{\theta}\right)\right].$ (5.58)
and
$-\frac{i}{2}\operatorname{Tr}\!\left[\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right),\sqrt{\rho_{\theta}}\right]H_{\theta}^{\star}\right]$
$=-\frac{1}{2}\operatorname{Tr}\!\left[\left(\sum_{r,s}\left(\frac{\lambda_{s}^{\frac{1}{2}}-\lambda_{r}^{\frac{1}{2}}}{\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}}\right)\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s}\right)\left(\sum_{\ell,m}\frac{\left(\lambda_{\ell}^{\frac{1}{2}}-\lambda_{m}^{\frac{1}{2}}\right)}{\left(\lambda_{\ell}+\lambda_{m}\right)\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)}\,\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\right)\right]$ (5.59)
$=-\frac{1}{2}\sum_{r,s,\ell,m}\frac{\left(\lambda_{s}^{\frac{1}{2}}-\lambda_{r}^{\frac{1}{2}}\right)\left(\lambda_{\ell}^{\frac{1}{2}}-\lambda_{m}^{\frac{1}{2}}\right)}{\left(\lambda_{r}^{\frac{1}{2}}+\lambda_{s}^{\frac{1}{2}}\right)\left(\lambda_{\ell}+\lambda_{m}\right)\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)}\operatorname{Tr}\!\left[\Pi_{r}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{s}\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\right]$ (5.60)
$=-\frac{1}{2}\sum_{\ell,m}\frac{\left(\lambda_{\ell}^{\frac{1}{2}}-\lambda_{m}^{\frac{1}{2}}\right)^{2}}{\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)^{2}\left(\lambda_{\ell}+\lambda_{m}\right)}\operatorname{Tr}\!\left[\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\left(\partial_{\theta}\rho_{\theta}\right)\right].$ (5.61)
Finally, we combine (5.55), (5.58), and (5.61) to conclude that
$\operatorname{Tr}\!\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right)^{2}\right]-\frac{i}{2}\operatorname{Tr}\!\left[\left[\left(\partial_{\theta}\sqrt{\rho_{\theta}}\right),\sqrt{\rho_{\theta}}\right]H_{\theta}^{\star}\right]$ (5.62)
$=\sum_{\ell,m}\left(\frac{1}{\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)^{2}}-\frac{\left(\lambda_{\ell}^{\frac{1}{2}}-\lambda_{m}^{\frac{1}{2}}\right)^{2}}{2\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)^{2}\left(\lambda_{\ell}+\lambda_{m}\right)}\right)\operatorname{Tr}\!\left[\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\left(\partial_{\theta}\rho_{\theta}\right)\right]$ (5.63)
$=\sum_{\ell,m}\frac{1}{\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)^{2}}\left(1-\frac{\left(\lambda_{\ell}^{\frac{1}{2}}-\lambda_{m}^{\frac{1}{2}}\right)^{2}}{2\left(\lambda_{\ell}+\lambda_{m}\right)}\right)\operatorname{Tr}\!\left[\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\left(\partial_{\theta}\rho_{\theta}\right)\right].$ (5.64)
Now consider that, for $x,y>0$,
$\frac{1}{\left(x^{\frac{1}{2}}+y^{\frac{1}{2}}\right)^{2}}\left(1-\frac{\left(x^{\frac{1}{2}}-y^{\frac{1}{2}}\right)^{2}}{2\left(x+y\right)}\right)$ $=\frac{1}{\left(x^{\frac{1}{2}}+y^{\frac{1}{2}}\right)^{2}}\left(\frac{2\left(x+y\right)-\left(x^{\frac{1}{2}}-y^{\frac{1}{2}}\right)^{2}}{2\left(x+y\right)}\right)$ (5.65)
$=\frac{1}{x+y+2\sqrt{xy}}\left(\frac{2\left(x+y\right)-\left(x+y-2\sqrt{xy}\right)}{2\left(x+y\right)}\right)$ (5.66)
$=\frac{1}{x+y+2\sqrt{xy}}\left(\frac{x+y+2\sqrt{xy}}{2\left(x+y\right)}\right)$ (5.67)
$=\frac{1}{2\left(x+y\right)},$ (5.68)
implying that
$\sum_{\ell,m}\left(\frac{1}{\left(\lambda_{\ell}^{\frac{1}{2}}+\lambda_{m}^{\frac{1}{2}}\right)^{2}}\left(1-\frac{\left(\lambda_{\ell}^{\frac{1}{2}}-\lambda_{m}^{\frac{1}{2}}\right)^{2}}{2\left(\lambda_{\ell}+\lambda_{m}\right)}\right)\right)\operatorname{Tr}\!\left[\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\left(\partial_{\theta}\rho_{\theta}\right)\right]$
$=\sum_{\ell,m}\frac{1}{2\left(\lambda_{\ell}+\lambda_{m}\right)}\operatorname{Tr}\!\left[\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\left(\partial_{\theta}\rho_{\theta}\right)\right]$ (5.69)
$=\frac{1}{4}\sum_{\ell,m}\frac{2}{\lambda_{\ell}+\lambda_{m}}\operatorname{Tr}\!\left[\Pi_{\ell}\left(\partial_{\theta}\rho_{\theta}\right)\Pi_{m}\left(\partial_{\theta}\rho_{\theta}\right)\right]$ (5.70)
$=\frac{1}{4}I_{F}\!\left(\theta;\left(\rho_{\theta}\right)_{\theta}\right),$ (5.71)
thus concluding the proof. ∎
5.2 SLD Fisher information of channels
The SLD Fisher information of a smooth family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of quantum channels is defined as follows [Fuj01]:
$I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}\right)$ $\coloneqq\sup_{\rho_{RA}\in\mathbb{D}}I_{F}\!\left(\theta;\left(\left(\operatorname{id}_{R}\otimes\mathcal{N}_{\theta}\right)(\rho_{RA})\right)_{\theta\in\Theta}\right).$ (5.72)
Let us suppose that the channel $\mathcal{N}_{\theta}$ has a Kraus decomposition as $\mathcal{N}_{\theta}(\rho)=\sum_{i}K_{i}^{\theta}\rho\left(K_{i}^{\theta}\right)^{{\dagger}}$, where $\left(K_{i}^{\theta}\right)_{\theta\in\Theta}$ is a smooth family of Kraus operators for each $i$, and let us further suppose that the number of Kraus operators for $\mathcal{N}_{\theta}$ is fixed for all $\theta\in\Theta$. Then we take a canonical isometric extension of the channel $\mathcal{N}_{\theta}$ as follows:
$V_{\theta}\coloneqq\sum_{i}|i\rangle\otimes K_{i}^{\theta},$ (5.73)
where $V_{\theta}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$.
The following theorem is strongly related to [FI08, Theorem 4], with a key difference being that the formulas for SLD Fisher information of channels are expressed below in terms of the smooth family $\left(V_{\theta}\right)_{\theta\in\Theta}$ of canonical isometric extensions of the channels in the family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$.
Theorem 13.
The SLD Fisher information of a smooth family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of quantum channels can be written as follows:
$\frac{1}{4}I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}\right)$ $=\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\partial_{\theta}\left[\left(U_{\theta}\otimes I_{RB}\right)V_{\theta}|\psi\rangle_{RA}\right]\right\|^{2}$ (5.74)
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\left(\partial_{\theta}V_{\theta}\right)-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta}\right\|^{2}$ (5.75)
$=\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\partial_{\theta}\left[\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right]\right\|^{2},$ (5.76)
where $V_{\theta}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ is a canonical isometric extension of $\mathcal{N}_{\theta}$, as defined in (5.73), and the optimizations in (5.75) and (5.76) are over every smooth family $\left(H_{\theta}\right)_{\theta\in\Theta}$ of Hermitian operators and every smooth family $\left(U_{\theta}\right)_{\theta\in\Theta}$ of unitary operators, respectively.
Proof.
Let us begin by noting that
$I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}\right)=\sup_{|\psi\rangle_{RA}\in\mathbb{P}}I_{F}\!\left(\theta;\left(\left(\operatorname{id}_{R}\otimes\mathcal{N}_{\theta}\right)(\psi_{RA})\right)_{\theta\in\Theta}\right),$ (5.77)
which follows from the data-processing inequality for the SLD Fisher information of states, purification, and the Schmidt decomposition theorem. Furthermore, in (5.77), the reference system $R$ and the channel input system $A$ are isomorphic. (See also [KW21b, Remark 10] in this context.) Consider that we can write an arbitrary state vector $|\psi\rangle_{RA}$ in terms of a unitary $U$ and a density operator $\rho$ as
$|\psi\rangle_{RA}=\left(U_{R}\otimes\sqrt{\rho}\right)|\Gamma\rangle.$ (5.78)
Then
$\left(U_{\theta}\otimes I_{RB}\right)V_{\theta}|\psi\rangle_{RA}=\sum_{i}U_{\theta}|i\rangle\otimes\left(U_{R}\otimes K_{i}^{\theta}\sqrt{\rho}\right)|\Gamma\rangle$ (5.79)
and, by appealing to Theorem 12, we can write
$\frac{1}{4}I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}\right)=$ $\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\partial_{\theta}\left[\left(U_{\theta}\otimes I_{RB}\right)V_{\theta}|\psi\rangle_{RA}\right]\right\|^{2}.$ (5.80)
This proves (5.74).
Now let us work with the objective function in (5.80). Defining
$|\psi^{\rho}\rangle_{RA}\coloneqq\left(I\otimes\sqrt{\rho}\right)|\Gamma\rangle,$ (5.81)
consider that
$\left\|\partial_{\theta}\left[\left(U_{\theta}\otimes I_{RB}\right)V_{\theta}|\psi\rangle_{RA}\right]\right\|^{2}$
$=\left\|\partial_{\theta}\left[\left(U_{\theta}\otimes I_{RB}\right)V_{\theta}|\psi^{\rho}\rangle_{RA}\right]\right\|^{2}$ (5.82)
$=\left\|\left[\left(\left(\partial_{\theta}U_{\theta}\right)\otimes I_{RB}\right)V_{\theta}+\left(U_{\theta}\otimes I_{RB}\right)\left(\partial_{\theta}V_{\theta}\right)\right]|\psi^{\rho}\rangle_{RA}\right\|^{2}$ (5.83)
$=\left\|U_{\theta}^{{\dagger}}\left[\left(\left(\partial_{\theta}U_{\theta}\right)\otimes I_{RB}\right)V_{\theta}+\left(U_{\theta}\otimes I_{RB}\right)\left(\partial_{\theta}V_{\theta}\right)\right]|\psi^{\rho}\rangle_{RA}\right\|^{2}$ (5.84)
$=\left\|\left[\left(U_{\theta}^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)\otimes I_{RB}\right)V_{\theta}+\left(U_{\theta}^{{\dagger}}U_{\theta}\otimes I_{RB}\right)\left(\partial_{\theta}V_{\theta}\right)\right]|\psi^{\rho}\rangle_{RA}\right\|^{2}$ (5.85)
$=\left\|\left[\partial_{\theta}V_{\theta}+\left(U_{\theta}^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)\otimes I_{RB}\right)V_{\theta}\right]|\psi^{\rho}\rangle_{RA}\right\|^{2}$ (5.86)
$=\left\|\left[\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{RB}\right)V_{\theta}\right]|\psi^{\rho}\rangle_{RA}\right\|^{2}$ (5.87)
$=\langle\psi^{\rho}|_{RA}\left[\partial_{\theta}V_{\theta}^{{\dagger}}+iV_{\theta}^{{\dagger}}\left(H_{\theta}\otimes I_{RB}\right)\right]\left[\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{RB}\right)V_{\theta}\right]|\psi^{\rho}\rangle_{RA}$ (5.88)
$=\operatorname{Tr}\!\left[\left[\partial_{\theta}V_{\theta}^{{\dagger}}+iV_{\theta}^{{\dagger}}\left(H_{\theta}\otimes I_{RB}\right)\right]\left[\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{RB}\right)V_{\theta}\right]\rho\right].$ (5.89)
To establish the equality in (5.87), we invoked the same reasoning around (5.16)–(5.20). Now observe that we have written the objective function in two different ways in (5.87) and (5.89) to see that it is is convex in $H_{\theta}$ and linear in $\rho$. Indeed, the function
$H_{\theta}\mapsto\left\|\left[\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{RB}\right)V_{\theta}\right]|\psi\rangle_{RA}\right\|^{2}$ (5.90)
is convex in $H_{\theta}$, being a composition of a convex function (the norm) and a convex and non-decreasing function (the square function $x\mapsto x^{2}$), and the function
$\rho\mapsto\operatorname{Tr}\!\left[\left[\partial_{\theta}V_{\theta}^{{\dagger}}+iV_{\theta}^{{\dagger}}\left(H_{\theta}\otimes I_{RB}\right)\right]\left[\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{RB}\right)V_{\theta}\right]\rho\right]$ (5.91)
is linear in $\rho$. Then we can apply the standard minimax theorem because the set of density operators is convex and compact and the set of Hermitian operators is convex. Thus,
$\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\partial_{\theta}\left[\left(U_{\theta}\otimes I_{RB}\right)V_{\theta}|\psi\rangle_{RA}\right]\right\|^{2}$
$=\sup_{\rho\in\mathbb{D}}\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\left[\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{RB}\right)V_{\theta}\right]|\psi^{\rho}\rangle_{RA}\right\|^{2}$ (5.92)
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\sup_{\rho\in\mathbb{D}}\left\|\left[\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{RB}\right)V_{\theta}\right]|\psi^{\rho}\rangle_{RA}\right\|^{2}$ (5.93)
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\left\|\left[\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{RB}\right)V_{\theta}\right]|\psi\rangle_{RA}\right\|^{2}$ (5.94)
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{RB}\right)V_{\theta}\right\|^{2}$ (5.95)
$=\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta}\right\|^{2}.$ (5.96)
This proves (5.75).
The final expression in (5.76) follows by noting that
$\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\|\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta}\right\|^{2}=\inf_{\left(U_{\theta}\right)_{\theta\in\Theta}}\left\|\partial_{\theta}\left[\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right]\right\|^{2},$ (5.97)
again as a consequence of the reasoning around (5.16)–(5.20). ∎
Remark 14.
Observe that the overlap term $\left\|\left(\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right)^{{\dagger}}\partial_{\theta}\left[\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right]\right\|$ simplifies as follows:
$\left\|\left(\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right)^{{\dagger}}\partial_{\theta}\left[\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right]\right\|$
$=\left\|V_{\theta}^{{\dagger}}\left(U_{\theta}^{{\dagger}}\otimes I_{B}\right)\partial_{\theta}\left[\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right]\right\|$ (5.98)
$=\left\|V_{\theta}^{{\dagger}}\left(U_{\theta}^{{\dagger}}\otimes I_{B}\right)\left[\left(\left(\partial_{\theta}U_{\theta}\right)\otimes I_{B}\right)V_{\theta}+\left(U_{\theta}\otimes I_{B}\right)\left(\partial_{\theta}V_{\theta}\right)\right]\right\|$ (5.99)
$=\left\|V_{\theta}^{{\dagger}}\left[\left(U_{\theta}^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)\otimes I_{B}\right)V_{\theta}+\left(\left(U_{\theta}\right)^{{\dagger}}\otimes I_{B}\right)\left(U_{\theta}\otimes I_{B}\right)\left(\partial_{\theta}V_{\theta}\right)\right]\right\|$ (5.100)
$=\left\|V_{\theta}^{{\dagger}}\left[\left(U_{\theta}^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)\otimes I_{B}\right)V_{\theta}+\partial_{\theta}V_{\theta}\right]\right\|$ (5.101)
$=\left\|V_{\theta}^{{\dagger}}\left(\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta}\right)\right\|,$ (5.102)
where again we define $H_{\theta}\coloneqq iU_{\theta}^{{\dagger}}\left(\partial_{\theta}U_{\theta}\right)$ (see (5.16)–(5.20)). The expression in (5.102) finds application in bounds on query complexity of channel estimation.
Remark 15.
Relevant for our applications in channel estimation, we can alternatively express the objective function in (5.75) and the overlap term in (5.102) in terms of Kraus operators $\left(K_{i}^{\theta}\right)_{i}$ of $\mathcal{N}_{\theta}$ as follows, by making use of (5.73). This connects to the notation used in [FI08, Section 4.1] and in several subsequent papers on channel estimation [DDKG12, KDD13, DDM14, ZJ21, KGADD23]. To begin with, consider that
$\left(U_{\theta}\otimes I_{B}\right)V_{\theta}$ $=\left(U_{\theta}\otimes I_{B}\right)\left(\sum_{j}|j\rangle\otimes K_{j}^{\theta}\right)$ (5.103)
$=\sum_{j}U_{\theta}|j\rangle\otimes K_{j}^{\theta}$ (5.104)
$=\sum_{j}\sum_{i}|i\rangle\!\langle i|U_{\theta}|j\rangle\otimes K_{j}^{\theta}$ (5.105)
$=\sum_{j}\sum_{i}|i\rangle\otimes u_{i,j}^{\theta}K_{j}^{\theta}$ (5.106)
$=\sum_{i}|i\rangle\otimes\sum_{j}u_{i,j}^{\theta}K_{j}^{\theta}$ (5.107)
$=\sum_{i}|i\rangle\otimes\widetilde{K}_{i}^{\theta},$ (5.108)
where we defined $u_{i,j}^{\theta}\coloneqq\langle i|U_{\theta}|j\rangle$ and the Kraus operators $\widetilde{K}_{i}^{\theta}\coloneqq\sum_{j}u_{i,j}^{\theta}K_{j}^{\theta}$. This implies that
$\partial_{\theta}\left[\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right]=\sum_{i}|i\rangle\otimes\partial_{\theta}\widetilde{K}_{i}^{\theta}.$ (5.109)
Then the objective function in (5.75) simplifies as follows:
$\left\|\partial_{\theta}\left[\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right]\right\|^{2}$ $=\left\|\sum_{i}|i\rangle\otimes\partial_{\theta}\widetilde{K}_{i}^{\theta}\right\|^{2}$ (5.110)
$=\left\|\left(\sum_{j}|j\rangle\otimes\partial_{\theta}\widetilde{K}_{j}^{\theta}\right)^{{\dagger}}\left(\sum_{i}|i\rangle\otimes\partial_{\theta}\widetilde{K}_{i}^{\theta}\right)\right\|$ (5.111)
$=\left\|\left(\sum_{j}\langle j|\otimes\left(\partial_{\theta}\widetilde{K}_{j}^{\theta}\right)^{{\dagger}}\right)\left(\sum_{i}|i\rangle\otimes\partial_{\theta}\widetilde{K}_{i}^{\theta}\right)\right\|$ (5.112)
$=\left\|\sum_{j}\left(\partial_{\theta}\widetilde{K}_{j}^{\theta}\right)^{{\dagger}}\partial_{\theta}\widetilde{K}_{j}^{\theta}\right\|,$ (5.113)
which is written in [FI08, Section 4.1] as $\left\|\alpha\right\|$, where
$\alpha\coloneqq\sum_{j}\left(\partial_{\theta}\widetilde{K}_{j}^{\theta}\right)^{{\dagger}}\partial_{\theta}\widetilde{K}_{j}^{\theta}.$ (5.114)
Additionally, the overlap term in (5.102) simplifies as follows:
$\left\|\left(\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right)^{{\dagger}}\partial_{\theta}\left[\left(U_{\theta}\otimes I_{B}\right)V_{\theta}\right]\right\|$ $=\left\|\left(\sum_{j}|j\rangle\otimes\widetilde{K}_{j}^{\theta}\right)^{{\dagger}}\left(\sum_{i}|i\rangle\otimes\partial_{\theta}\widetilde{K}_{i}^{\theta}\right)\right\|$ (5.115)
$=\left\|\left(\sum_{j}\langle j|\otimes\left(\widetilde{K}_{j}^{\theta}\right)^{{\dagger}}\right)\left(\sum_{i}|i\rangle\otimes\partial_{\theta}\widetilde{K}_{i}^{\theta}\right)\right\|$ (5.116)
$=\left\|\sum_{j}\left(\widetilde{K}_{j}^{\theta}\right)^{{\dagger}}\partial_{\theta}\widetilde{K}_{j}^{\theta}\right\|,$ (5.117)
which is written in [FI08, Section 4.1] as $\left\|\beta\right\|$, where
$\beta\coloneqq\sum_{j}\left(\widetilde{K}_{j}^{\theta}\right)^{{\dagger}}\partial_{\theta}\widetilde{K}_{j}^{\theta}.$ (5.118)
The following theorem was stated in [YF17, Eq. (18)]; below we provide a proof for it.
Theorem 16.
Let $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ be a smooth family of quantum channels. The following equalities hold:
$d_{B}^{2}\!\left(\mathcal{N}_{\theta},\mathcal{N}_{\theta+\delta}\right)$ $=\frac{\delta^{2}}{4}I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}\right)_{\theta}\right)+o(\delta^{2}),$ (5.119)
$I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}\right)_{\theta}\right)$ $=2\left.\frac{\partial^{2}}{\partial\delta^{2}}d_{B}^{2}(\mathcal{N}_{\theta},\mathcal{N}_{\theta+\delta})\right|_{\delta=0}$ (5.120)
$=4\lim_{\delta\to 0}\frac{d_{B}^{2}(\mathcal{N}_{\theta},\mathcal{N}_{\theta+\delta})}{\delta^{2}}.$ (5.121)
Proof.
Recall (5.77). For a fixed pure input state $\psi_{RA}$, applying (5.3) leads to
$d_{B}^{2}\!\left(\mathcal{N}_{\theta}\!\left(\psi_{RA}\right),\mathcal{N}_{\theta+\delta}\!\left(\psi_{RA}\right)\right)$ $=\frac{\delta^{2}}{4}I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}\!\left(\psi_{RA}\right)\right)_{\theta}\right)+o(\delta^{2}).$ (5.122)
The convergence in the $o(\delta^{2})$ term is uniform in $\psi_{RA}$, as a consequence of the following general fact: Let $K$ be compact and let $f\colon K\times\left[-\varepsilon,\varepsilon\right]\to\mathbb{R}$ be such that $f(\psi,\cdot)$ is $C^{2}$ for each $\psi\in K$ and $\left(\psi,\delta\right)\mapsto f^{\prime\prime}(\psi,\delta)$ is continuous. Then
$f(\psi,\delta)=f(\psi,0)+f^{\prime}(\psi,0)\delta+\frac{1}{2}f^{\prime\prime}(\psi,0)\delta^{2}+o(\delta^{2})$ (5.123)
uniformly in $\psi\in K$. Applying this statement to our case, the equality in (5.122) holds because the set of pure states on $RA$ is compact. Then we conclude that
$\sup_{\psi_{RA}\in\mathbb{P}}d_{B}^{2}\!\left(\mathcal{N}_{\theta}\!\left(\psi_{RA}\right),\mathcal{N}_{\theta+\delta}\!\left(\psi_{RA}\right)\right)=\frac{\delta^{2}}{4}\left[\sup_{\psi_{RA}\in\mathbb{P}}I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}\!\left(\psi_{RA}\right)\right)_{\theta}\right)\right]+o(\delta^{2}),$ (5.124)
given that the convergence is uniform and the supremum is taken over a compact set of states, so that the maximizing $\psi_{RA}$ is smooth in $\delta$ near 0. This equality is equivalent to
$d_{B}^{2}\!\left(\mathcal{N}_{\theta},\mathcal{N}_{\theta+\delta}\right)=\frac{\delta^{2}}{4}I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}\right)_{\theta}\right)+o(\delta^{2}).$ (5.125)
Then (5.120) and (5.121) are immediate consequences of (5.125). ∎
6 Upper bounds on Bures distance of channels
In this section, we present upper bounds on the Bures distance of channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ in both the parallel and adaptive settings. The bounds are expressed in terms of a canonical isometric extension $V_{i}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ of the channel $\mathcal{N}_{i}$, for $i\in\left\{1,2\right\}$, as given in (4.38) and (4.39).
6.1 Upper bound on Bures distance of channels in the parallel setting
Ref. [YF17, Eq. (23)] established an upper bound on the squared Bures distance of the tensor-power channels $\mathcal{N}_{1}^{\otimes n}$ and $\mathcal{N}_{2}^{\otimes n}$. Here we restate this upper bound in a different form, in terms of canonical isometric extensions of the channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, and we provide an alternate proof based on basic properties of isometries and the spectral norm.
Theorem 17.
For channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ with canonical isometric extensions $V_{1}$ and $V_{2}$, the following inequality holds for all $n\in\mathbb{N}$:
$d_{B}^{2}\!\left(\mathcal{N}_{1}^{\otimes n},\mathcal{N}_{2}^{\otimes n}\right)\leq n\inf_{W\in\mathbb{B}}\left\{2\left\|I-\Re\!\left[M_{W}\right]\right\|+\left(n-1\right)\left\|I-M_{W}\right\|^{2}\right\},$ (6.1)
where
$M_{W}\equiv V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}.$ (6.2)
Proof.
Let $V_{i}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ be a canonical isometric extension of $\mathcal{N}_{i}$, and set
$\hat{V}_{1}$ $\equiv|0\rangle\otimes V_{1},$ (6.3)
$\hat{V}_{2}$ $\equiv\left(U_{W}\otimes I\right)|0\rangle\otimes V_{2},$ (6.4)
where $|0\rangle\in\mathbb{C}^{2}$ and $U_{W}$ is an arbitrary unitary acting nontrivially on the first two systems of $\mathbb{C}^{2}\otimes\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$. As such, we can apply Theorem 10 to $d_{B}^{2}\!\left(\mathcal{N}_{1}^{\otimes n},\mathcal{N}_{2}^{\otimes n}\right)$ to conclude that
$d_{B}^{2}\!\left(\mathcal{N}_{1}^{\otimes n},\mathcal{N}_{2}^{\otimes n}\right)\leq\left\|\hat{V}_{1}^{\otimes n}-\hat{V}_{2}^{\otimes n}\right\|^{2}.$ (6.5)
Continuing, we find that
$\left\|\hat{V}_{1}^{\otimes n}-\hat{V}_{2}^{\otimes n}\right\|^{2}$
$=\left\|\sum_{i=1}^{n}\hat{V}_{1}^{\otimes i-1}\otimes\left[\hat{V}_{1}-\hat{V}_{2}\right]_{i}\otimes\hat{V}_{2}^{\otimes n-i}\right\|^{2}$ (6.6)
$=\left\|\left(\sum_{j=1}^{n}\hat{V}_{1}^{\otimes j-1}\otimes\left[\hat{V}_{1}-\hat{V}_{2}\right]_{j}\otimes\hat{V}_{2}^{\otimes n-j}\right)^{{\dagger}}\left(\sum_{i=1}^{n}\hat{V}_{1}^{\otimes i-1}\otimes\left[\hat{V}_{1}-\hat{V}_{2}\right]_{i}\otimes\hat{V}_{2}^{\otimes n-i}\right)\right\|$ (6.7)
$=\left\|\left(\sum_{j=1}^{n}\left(\hat{V}_{1}^{{\dagger}}\right)^{\otimes j-1}\otimes\left[\hat{V}_{1}^{{\dagger}}-\hat{V}_{2}^{{\dagger}}\right]_{j}\otimes\left(\hat{V}_{2}^{{\dagger}}\right)^{\otimes n-j}\right)\left(\sum_{i=1}^{n}\hat{V}_{1}^{\otimes i-1}\otimes\left[\hat{V}_{1}-\hat{V}_{2}\right]_{i}\otimes\hat{V}_{2}^{\otimes n-i}\right)\right\|$ (6.8)
$=\left\|\sum_{i,j=1}^{n}\left(\left(\hat{V}_{1}^{{\dagger}}\right)^{\otimes j-1}\otimes\left[\hat{V}_{1}^{{\dagger}}-\hat{V}_{2}^{{\dagger}}\right]_{j}\otimes\left(\hat{V}_{2}^{{\dagger}}\right)^{\otimes n-j}\right)\left(\hat{V}_{1}^{\otimes i-1}\otimes\left[\hat{V}_{1}-\hat{V}_{2}\right]_{i}\otimes\hat{V}_{2}^{\otimes n-i}\right)\right\|$ (6.9)
$\leq\sum_{i=1}^{n}\left\|\left(\left(\hat{V}_{1}^{{\dagger}}\right)^{\otimes i-1}\otimes\left[\hat{V}_{1}^{{\dagger}}-\hat{V}_{2}^{{\dagger}}\right]_{i}\otimes\left(\hat{V}_{2}^{{\dagger}}\right)^{\otimes n-i}\right)\left(\hat{V}_{1}^{\otimes i-1}\otimes\left[\hat{V}_{1}-\hat{V}_{2}\right]_{i}\otimes\hat{V}_{2}^{\otimes n-i}\right)\right\|$
$\qquad+\sum_{i,j=1:i<j}^{n}\left\|\left(\left(\hat{V}_{1}^{{\dagger}}\right)^{\otimes j-1}\otimes\left[\hat{V}_{1}^{{\dagger}}-\hat{V}_{2}^{{\dagger}}\right]_{j}\otimes\left(\hat{V}_{2}^{{\dagger}}\right)^{\otimes n-j}\right)\left(\hat{V}_{1}^{\otimes i-1}\otimes\left[\hat{V}_{1}-\hat{V}_{2}\right]_{i}\otimes\hat{V}_{2}^{\otimes n-i}\right)\right\|$
$\qquad+\sum_{i,j=1:i>j}^{n}\left\|\left(\left(\hat{V}_{1}^{{\dagger}}\right)^{\otimes j-1}\otimes\left[\hat{V}_{1}^{{\dagger}}-\hat{V}_{2}^{{\dagger}}\right]_{j}\otimes\left(\hat{V}_{2}^{{\dagger}}\right)^{\otimes n-j}\right)\left(\hat{V}_{1}^{\otimes i-1}\otimes\left[\hat{V}_{1}-\hat{V}_{2}\right]_{i}\otimes\hat{V}_{2}^{\otimes n-i}\right)\right\|$ (6.10)
$=\sum_{i=1}^{n}\left\|\left[\hat{V}_{1}^{{\dagger}}-\hat{V}_{2}^{{\dagger}}\right]\left[\hat{V}_{1}-\hat{V}_{2}\right]\right\|+\sum_{i,j=1:i<j}^{n}\left\|\left[\hat{V}_{1}^{{\dagger}}\right]_{i}\left[\hat{V}_{1}-\hat{V}_{2}\right]_{i}\right\|\left\|\left[\hat{V}_{1}^{{\dagger}}-\hat{V}_{2}^{{\dagger}}\right]_{j}\left[\hat{V}_{2}\right]_{j}\right\|$
$\qquad+\sum_{i,j=1:i>j}^{n}\left\|\left[\hat{V}_{1}^{{\dagger}}-\hat{V}_{2}^{{\dagger}}\right]_{j}\left[\hat{V}_{1}\right]_{j}\right\|\left\|\left[\hat{V}_{2}^{{\dagger}}\right]_{i}\left[\hat{V}_{1}-\hat{V}_{2}\right]_{i}\right\|$ (6.11)
$=n\left\|\hat{V}_{1}-\hat{V}_{2}\right\|^{2}+\sum_{i,j=1:i<j}^{n}\left\|\hat{V}_{1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)\right\|\left\|\left(\hat{V}_{1}^{{\dagger}}-\hat{V}_{2}^{{\dagger}}\right)\hat{V}_{2}\right\|$
$\qquad+\sum_{i,j=1:i>j}^{n}\left\|\left(\hat{V}_{1}^{{\dagger}}-\hat{V}_{2}^{{\dagger}}\right)\hat{V}_{1}\right\|\left\|\hat{V}_{2}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)\right\|$ (6.12)
$=n\left\|\hat{V}_{1}-\hat{V}_{2}\right\|^{2}+\frac{n\left(n-1\right)}{2}\left\|I-\hat{V}_{1}^{{\dagger}}\hat{V}_{2}\right\|\left\|\hat{V}_{1}^{{\dagger}}\hat{V}_{2}-I\right\|+\frac{n\left(n-1\right)}{2}\left\|I-\hat{V}_{2}^{{\dagger}}\hat{V}_{1}\right\|\left\|\hat{V}_{2}^{{\dagger}}\hat{V}_{1}-I\right\|$ (6.13)
$=n\left\|\hat{V}_{1}-\hat{V}_{2}\right\|^{2}+n\left(n-1\right)\left\|I-\hat{V}_{1}^{{\dagger}}\hat{V}_{2}\right\|^{2}$ (6.14)
$=n\left(2\left\|I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right\|+\left(n-1\right)\left\|I-V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right\|^{2}\right).$ (6.15)
The last equality follows by expanding the first norm as
$\left\|\hat{V}_{1}-\hat{V}_{2}\right\|^{2}$ $=\left\|\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)\right\|$ (6.16)
$=\left\|2I-\hat{V}_{2}^{{\dagger}}\hat{V}_{1}-\hat{V}_{1}^{{\dagger}}\hat{V}_{2}\right\|$ (6.17)
$=2\left\|I-\Re\!\left[\hat{V}_{1}^{{\dagger}}\hat{V}_{2}\right]\right\|,$ (6.18)
and observing that
$\hat{V}_{1}^{{\dagger}}\hat{V}_{2}$ $=\left(\langle 0|\otimes V_{1}^{{\dagger}}\right)\left(U_{W}\otimes I\right)\left(|0\rangle\otimes V_{2}\right)$ (6.19)
$=V_{1}^{{\dagger}}\left[\left(\langle 0|\otimes I\right)U_{W}\left(|0\rangle\otimes I\right)\otimes I_{B}\right]V_{2}$ (6.20)
$=V_{1}^{{\dagger}}\left[W_{E}\otimes I_{B}\right]V_{2}.$ (6.21)
Given that $W_{E}$ is an arbitrary contraction, we can obtain the tightest upper bound using this approach by minimizing over all such contractions, thus completing the proof. ∎
Remark 18.
To see that the upper bound in (6.1) is equivalent to [YF17, Eq. (23)], we appeal to Remark 11 and define $K_{W}\coloneqq\sum_{i,j}W_{i,j}K_{i}^{{\dagger}}L_{j}$. Then
$2\left\|I-\Re\!\left[M_{W}\right]\right\|+\left(n-1\right)\left\|I-M_{W}\right\|^{2}=\left\|2I-K_{W}-K_{W}^{{\dagger}}\right\|+\left(n-1\right)\left\|I-K_{W}\right\|^{2},$ (6.22)
thus establishing the equivalence.
6.2 Upper bound on Bures distance of channels in the adaptive setting
For channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, a general adaptive $n$-query protocol consists of a tuple $\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}$ of channels. Applying the protocol to $n$ queries of the channel $\mathcal{N}_{i}$ leads to the following channel $\mathcal{P}_{i}^{(n)}$, for $i\in\left\{1,2\right\}$:
$\mathcal{P}_{i}^{(n)}\coloneqq\left(\operatorname{id}\otimes\mathcal{N}_{i}\right)\circ\mathcal{A}_{n-1}\circ\cdots\circ\mathcal{A}_{2}\circ\left(\operatorname{id}\otimes\mathcal{N}_{i}\right)\circ\mathcal{A}_{1}\circ\left(\operatorname{id}\otimes\mathcal{N}_{i}\right).$ (6.23)
This in turn leads to the following distinguishability measure for channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$:
$\sup_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}d_{B}^{2}\!\left(\mathcal{P}_{1}^{(n)},\mathcal{P}_{2}^{(n)}\right),$ (6.24)
which is related to the strategy fidelity of [GRS18] and is a special case of the generalized strategy divergence introduced in [WW19, Definition 1].
In Theorem 19 below, we establish an upper bound on the $n$-query adaptive squared Bures distance of channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$:
Theorem 19.
For channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ with canonical isometric extensions $V_{1}$ and $V_{2}$, the following inequality holds for all $n\in\mathbb{N}$:
$\sup_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}d_{B}^{2}\!\left(\mathcal{P}_{1}^{(n)},\mathcal{P}_{2}^{(n)}\right)\leq\\ n\inf_{W\in\mathbb{B}}\left\{2\left\|I-\Re\!\left[M_{W}\right]\right\|+\left(n-1\right)\left(2\left\|I-\Re\!\left[M_{W}\right]\right\|\right)^{\frac{1}{2}}\left\|I-M_{W}\right\|\right\},$ (6.25)
where
$M_{W}\equiv V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}.$ (6.26)
Proof.
Let us begin by observing that the channel $\mathcal{P}_{i}^{(n)}$ in (6.23) has an isometric extension of the following form:
$\left(I\otimes\hat{V}_{i}\right)A_{n-1}\cdots A_{2}\left(I\otimes\hat{V}_{i}\right)A_{1}\left(I\otimes\hat{V}_{i}\right),$ (6.27)
where $\hat{V}_{i}$ is chosen as in (6.3)–(6.4) and $A_{i}$ is an isometric extension of $\mathcal{A}_{i}$. Suppressing implicit tensor products with identity in what follows, we can apply Theorem 10 to $d_{B}^{2}\!\left(\mathcal{P}_{i}^{(n)},\mathcal{P}_{2}^{(n)}\right)$ to conclude that
$d_{B}^{2}\!\left(\mathcal{P}_{i}^{(n)},\mathcal{P}_{2}^{(n)}\right)\leq\left\|\hat{V}_{1}A_{n-1}\cdots A_{2}\hat{V}_{1}A_{1}\hat{V}_{1}-\hat{V}_{2}A_{n-1}\cdots A_{2}\hat{V}_{2}A_{1}\hat{V}_{2}\right\|^{2}.$ (6.28)
Continuing, consider that
$\left\|\hat{V}_{1}A_{n-1}\cdots A_{2}\hat{V}_{1}A_{1}\hat{V}_{1}-\hat{V}_{2}A_{n-1}\cdots A_{2}\hat{V}_{2}A_{1}\hat{V}_{2}\right\|^{2}$
$=\left\|\sum_{i=1}^{n}\hat{V}_{1}A_{n-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right\|^{2}$ (6.29)
$=\left\|\begin{array}[]{c}\left(\sum_{j=1}^{n}\hat{V}_{1}A_{n-1}\cdots A_{j+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{j-1}\cdots A_{1}\hat{V}_{2}\right)^{{\dagger}}\times\\ \left(\sum_{i=1}^{n}\hat{V}_{1}A_{n-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right)\end{array}\right\|$ (6.32)
$=\left\|\begin{array}[]{c}\left(\sum_{j=1}^{n}\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{1}^{{\dagger}}\right)\times\\ \left(\sum_{i=1}^{n}\hat{V}_{1}A_{n-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right)\end{array}\right\|$ (6.35)
$=\left\|\begin{array}[]{c}\sum_{i,j=1}^{n}\left(\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{1}^{{\dagger}}\right)\times\\ \left(\hat{V}_{1}A_{n-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right)\end{array}\right\|$ (6.38)
$\leq\sum_{i=1}^{n}\left\|\begin{array}[]{c}\left(\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{i-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}A_{i+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{1}^{{\dagger}}\right)\times\\ \left(\hat{V}_{1}A_{n-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right)\end{array}\right\|$ (6.41)
$\quad+\sum_{i,j=1:i>j}^{n}\left\|\begin{array}[]{c}\left(\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{1}^{{\dagger}}\right)\times\\ \left(\hat{V}_{1}A_{n-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right)\end{array}\right\|$ (6.44)
$\quad+\sum_{i,j=1:i<j}^{n}\left\|\begin{array}[]{c}\left(\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{1}^{{\dagger}}\right)\times\\ \left(\hat{V}_{1}A_{n-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right)\end{array}\right\|.$ (6.47)
Let us bound the $i$th term of the first sum in (6.47) as follows:
$\left\|\left(\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{i-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}A_{i+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{1}^{{\dagger}}\right)\left(\hat{V}_{1}A_{n-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right)\right\|$
$=\left\|\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{i-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right\|$ (6.48)
$\leq\left\|\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)\right\|\left\|\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{i-1}^{{\dagger}}A_{i-1}\cdots A_{1}\hat{V}_{2}\right\|$ (6.49)
$=\left\|\hat{V}_{1}-\hat{V}_{2}\right\|^{2}.$ (6.50)
Also, for $i>j$, let us bound the $(i,j)$th term of the second sum in (6.47) as follows:
$\left\|\left(\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{1}^{{\dagger}}\right)\left(\hat{V}_{1}A_{n-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right)\right\|$
$=\left\|\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}\hat{V}_{1}A_{j-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right\|$ (6.51)
$\leq\left\|\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\right\|\left\|\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}\hat{V}_{1}\right\|\left\|A_{j-1}\cdots A_{i+1}\right\|\left\|\hat{V}_{1}-\hat{V}_{2}\right\|\left\|A_{i-1}\cdots A_{1}\hat{V}_{2}\right\|$ (6.52)
$=\left\|\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}\hat{V}_{1}\right\|\left\|\hat{V}_{1}-\hat{V}_{2}\right\|$ (6.53)
$=\left\|I-\hat{V}_{2}^{{\dagger}}\hat{V}_{1}\right\|\left\|\hat{V}_{1}-\hat{V}_{2}\right\|$ (6.54)
$=\left\|\hat{V}_{1}-\hat{V}_{2}\right\|\left\|I-\hat{V}_{1}^{{\dagger}}\hat{V}_{2}\right\|.$ (6.55)
For $i<j$, let us bound the $(i,j)$th term of the third sum in (6.47) as follows:
$\left\|\left(\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{1}^{{\dagger}}\right)\left(\hat{V}_{1}A_{n-1}\cdots A_{i+1}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right)\right\|$
$=\left\|\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}A_{j+1}^{{\dagger}}\cdots A_{i-1}^{{\dagger}}\hat{V}_{1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)A_{i-1}\cdots A_{1}\hat{V}_{2}\right\|$ (6.56)
$\leq\left\|\hat{V}_{2}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\right\|\left\|\left(\hat{V}_{1}-\hat{V}_{2}\right)^{{\dagger}}\right\|\left\|A_{j+1}^{{\dagger}}\cdots A_{i-1}^{{\dagger}}\right\|\left\|\hat{V}_{1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)\right\|\left\|A_{i-1}\cdots A_{1}\hat{V}_{2}\right\|$ (6.57)
$=\left\|\hat{V}_{1}-\hat{V}_{2}\right\|\left\|\hat{V}_{1}^{{\dagger}}\left(\hat{V}_{1}-\hat{V}_{2}\right)\right\|$ (6.58)
$=\left\|\hat{V}_{1}-\hat{V}_{2}\right\|\left\|I-\hat{V}_{1}^{{\dagger}}\hat{V}_{2}\right\|.$ (6.59)
Then, applying (6.28), (6.47), (6.50), (6.55), and (6.59), we conclude that
$d_{B}^{2}\!\left(\mathcal{P}_{i}^{(n)},\mathcal{P}_{2}^{(n)}\right)$
$\leq n\left\|\hat{V}_{1}-\hat{V}_{2}\right\|^{2}+\frac{n\left(n-1\right)}{2}\left\|\hat{V}_{1}-\hat{V}_{2}\right\|\left\|I-\hat{V}_{1}^{{\dagger}}\hat{V}_{2}\right\|+\frac{n\left(n-1\right)}{2}\left\|\hat{V}_{1}-\hat{V}_{2}\right\|\left\|I-\hat{V}_{1}^{{\dagger}}\hat{V}_{2}\right\|$ (6.60)
$=n\left\|\hat{V}_{1}-\hat{V}_{2}\right\|^{2}+n\left(n-1\right)\left\|\hat{V}_{1}-\hat{V}_{2}\right\|\left\|I-\hat{V}_{1}^{{\dagger}}\hat{V}_{2}\right\|,$ (6.61)
thus concluding the proof after applying (6.16)–(6.21) and the reasoning thereafter. ∎
Remark 20.
We can alternatively express (6.25) using the notation of [YF17]. Appealing to Remark 11 and defining $K_{W}\coloneqq\sum_{i,j}W_{i,j}K_{i}^{{\dagger}}L_{j}$, we can write the objective function in (6.25) as follows:
$2\left\|I-\Re\!\left[M_{W}\right]\right\|+\left(n-1\right)\left(2\left\|I-\Re\!\left[M_{W}\right]\right\|\right)^{\frac{1}{2}}\left\|I-M_{W}\right\|\\ =\left\|2I-K_{W}-K_{W}^{{\dagger}}\right\|+\left(n-1\right)\left\|2I-K_{W}-K_{W}^{{\dagger}}\right\|^{\frac{1}{2}}\left\|I-K_{W}\right\|.$ (6.62)
Remark 21.
The upper bound in Theorem 17 does not exceed the upper bound in Theorem 19, consistent with the intuition that the Bures distance for parallel protocols should not exceed the Bures distance of adaptive protocols, given that parallel protocols are contained with the set of all adaptive protocols. This can be seen by comparing the objective functions in Theorems 17 and 19 and noting that the following inequality holds for every contraction $W$:
$\left\|I-M_{W}\right\|$ $=\left\|\left(I-M_{W}\right)^{{\dagger}}\left(I-M_{W}\right)\right\|^{\frac{1}{2}}$ (6.63)
$=\left\|I-M_{W}-M_{W}^{{\dagger}}+M_{W}^{{\dagger}}M_{W}\right\|^{\frac{1}{2}}$ (6.64)
$\leq\left\|I-M_{W}-M_{W}^{{\dagger}}+I\right\|^{\frac{1}{2}}$ (6.65)
$=\left(2\left\|I-\Re\!\left[M_{W}\right]\right\|\right)^{\frac{1}{2}},$ (6.66)
where the inequality follows because
$I-M_{W}-M_{W}^{{\dagger}}+M_{W}^{{\dagger}}M_{W}$ $\leq I-M_{W}-M_{W}^{{\dagger}}+I$ (6.67)
$\Longleftrightarrow\qquad M_{W}^{{\dagger}}M_{W}$ $\leq I,$ (6.68)
with the final inequality in (6.68) holding because $M_{W}$ is a contraction, itself being a composition of contractions.
7 Upper bounds on SLD Fisher information of channels
In this section, we present upper bounds on the SLD Fisher information of a smooth family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of channels, in both the parallel and adaptive settings. The bounds are expressed in terms of a canonical isometric extension $V_{\theta}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ of the channel $\mathcal{N}_{\theta}$, as given in (5.73).
7.1 Upper bound on SLD Fisher information in the parallel setting
[FI08, Section 4.1] established an upper bound on the SLD Fisher information of the smooth family $\smash{\left(\mathcal{N}_{\theta}^{\otimes n}\right)_{\theta\in\Theta}}$ of channels, for all $n\in\mathbb{N}$. This upper bound has subsequently been employed in several foundational papers on quantum estimation theory [DDKG12, KDD13, DDM14, ZJ21]. Here we restate this upper bound in a different form, in terms of canonical isometric extensions of the channel family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$, and we provide an alternate proof based on basic properties of isometries and the spectral norm.
Theorem 22.
For a smooth family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of quantum channels, the following inequality holds for all $n\in\mathbb{N}$:
$\frac{1}{4}I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}^{\otimes n}\right)_{\theta\in\Theta}\right)$ $\leq n\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{\left\|M_{H_{\theta}}\right\|^{2}+\left(n-1\right)\left\|V_{\theta}^{{\dagger}}M_{H_{\theta}}\right\|^{2}\right\},$ (7.1)
where $V_{\theta}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ is a canonical isometric extension of $\mathcal{N}_{\theta}$ of the form in (5.73), the optimization is over every smooth family $\left(H_{\theta}\right)_{\theta\in\Theta}$ of Hermitian operators, and
$M_{H_{\theta}}\equiv\left(\partial_{\theta}V_{\theta}\right)-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta}.$ (7.2)
Proof.
Let us employ the shorthand
$\hat{V}_{\theta}\equiv\left(U_{\theta}\otimes I_{B}\right)V_{\theta},$ (7.3)
where $\left(U_{\theta}\right)_{\theta\in\Theta}$ is a smooth family of unitary operators, so that
$\partial_{\theta}\hat{V}_{\theta}=\partial_{\theta}\left[\left(U_{\theta}\otimes I_{RB}\right)V_{\theta}\right]=\left(\partial_{\theta}V_{\theta}\right)-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta},$ (7.4)
where $\left(H_{\theta}\right)_{\theta\in\Theta}$ is a smooth family of Hermitian operators, which follows from reasoning similar to that given for (5.87). We can use the fact that $\hat{V}_{\theta}^{\otimes n}$ is a particular isometric extension of $\mathcal{N}_{\theta}^{\otimes n}$ and apply Theorem 13 to conclude that
$\frac{1}{4}I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}^{\otimes n}\right)_{\theta\in\Theta}\right)$
$\leq\left\|\partial_{\theta}\left[\hat{V}_{\theta}^{\otimes n}\right]\right\|^{2}$ (7.5)
$=\left\|\sum_{i=1}^{n}\hat{V}_{\theta}^{\otimes i-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}\right]\otimes\hat{V}_{\theta}^{\otimes n-i}\right\|^{2}$ (7.6)
$=\left\|\left(\sum_{j=1}^{n}\hat{V}_{\theta}^{\otimes j-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}\right]_{j}\otimes\hat{V}_{\theta}^{\otimes n-j}\right)^{{\dagger}}\left(\sum_{i=1}^{n}\hat{V}_{\theta}^{\otimes i-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}\right]_{i}\otimes\hat{V}_{\theta}^{\otimes n-i}\right)\right\|$ (7.7)
$=\left\|\left(\sum_{j=1}^{n}\left(\hat{V}_{\theta}^{{\dagger}}\right)^{\otimes j-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right]_{j}\otimes\left(\hat{V}_{\theta}^{{\dagger}}\right)^{\otimes n-j}\right)\left(\sum_{i=1}^{n}\hat{V}_{\theta}^{\otimes i-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}\right]_{i}\otimes\hat{V}_{\theta}^{\otimes n-i}\right)\right\|$ (7.8)
$=\left\|\sum_{i,j=1}^{n}\left(\left(\hat{V}_{\theta}^{{\dagger}}\right)^{\otimes j-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right]_{j}\otimes\left(\hat{V}_{\theta}^{{\dagger}}\right)^{\otimes n-j}\right)\left(\hat{V}_{\theta}^{\otimes i-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}\right]_{i}\otimes\hat{V}_{\theta}^{\otimes n-i}\right)\right\|$ (7.9)
$\leq\sum_{i=1}^{n}\left\|\left(\left(\hat{V}_{\theta}^{{\dagger}}\right)^{\otimes i-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right]_{i}\otimes\left(\hat{V}_{\theta}^{{\dagger}}\right)^{\otimes n-i}\right)\left(\hat{V}_{\theta}^{\otimes i-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}\right]_{i}\otimes\hat{V}_{\theta}^{\otimes n-i}\right)\right\|$
$\qquad+\sum_{i,j=1:i\neq j}^{n}\left\|\left(\left(\hat{V}_{\theta}^{{\dagger}}\right)^{\otimes j-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right]_{j}\otimes\left(\hat{V}_{\theta}^{{\dagger}}\right)^{\otimes n-j}\right)\left(\hat{V}_{\theta}^{\otimes i-1}\otimes\left[\partial_{\theta}\hat{V}_{\theta}\right]_{i}\otimes\hat{V}_{\theta}^{\otimes n-i}\right)\right\|$ (7.10)
$=\sum_{i=1}^{n}\left\|\left(\hat{V}_{\theta}^{{\dagger}}\hat{V}_{\theta}\right)^{\otimes i-1}\otimes\left[\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)\left(\partial_{\theta}\hat{V}_{\theta}\right)\right]_{i}\otimes\left(\hat{V}_{\theta}^{{\dagger}}\hat{V}_{\theta}\right)^{\otimes n-i}\right\|$
$\qquad+\sum_{i,j=1:i\neq j}^{n}\left\|\hat{V}_{\theta}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}\right)\right\|\left\|\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)\hat{V}_{\theta}\right\|$ (7.11)
$=\sum_{i=1}^{n}\left\|\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)\left(\partial_{\theta}\hat{V}_{\theta}\right)\right\|+\sum_{i,j=1:i\neq j}^{n}\left\|\hat{V}_{\theta}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}\right)\right\|^{2}$ (7.12)
$=\sum_{i=1}^{n}\left\|\partial_{\theta}\hat{V}_{\theta}\right\|^{2}+\sum_{i,j=1:i\neq j}^{n}\left\|\hat{V}_{\theta}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}\right)\right\|^{2}$ (7.13)
$=n\left\|\partial_{\theta}\hat{V}_{\theta}\right\|^{2}+n\left(n-1\right)\left\|\hat{V}_{\theta}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}\right)\right\|^{2}.$ (7.14)
The first inequality follows from the triangle inequality for the spectral norm. The other equalities follow from the spectral norm being multiplicative under tensor products and it being equal to one for the identity operator. Applying (7.4) and Remark 14 and minimizing over every smooth family $\left(H_{\theta}\right)_{\theta\in\Theta}$ of Hermitian operators, we conclude the proof. ∎
7.2 Upper bound on SLD Fisher information in the adaptive setting
For a smooth family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of quantum channels, a general adaptive $n$-query protocol for channel estimation consists of a tuple $\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}$ of channels. Applying the protocol to $n$ queries of the channel $\mathcal{N}_{\theta}$ leads to the following channel $\mathcal{P}_{\theta}^{(n)}$ for $\theta\in\Theta$:
$\mathcal{P}_{\theta}^{(n)}\coloneqq\left(\operatorname{id}\otimes\mathcal{N}_{\theta}\right)\circ\mathcal{A}_{n-1}\circ\cdots\circ\mathcal{A}_{2}\circ\left(\operatorname{id}\otimes\mathcal{N}_{\theta}\right)\circ\mathcal{A}_{1}\circ\left(\operatorname{id}\otimes\mathcal{N}_{\theta}\right).$ (7.15)
This in turn leads to the following information measure for the family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$:
$\sup_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}I_{F}\!\left(\theta;\left(\mathcal{P}_{\theta}^{(n)}\right)_{\theta\in\mathbb{R}}\right),$ (7.16)
which is related to the general measure considered in [KW21b, Theorem 18].
Ref. [KGADD23, Eq. (11)] established an upper bound on the $n$-query adaptive SLD Fisher information in (7.16), for all $n\in\mathbb{N}$, which improves upon a previous bound from [DDM14, Eq. (9)]. Here we restate this upper bound in a different form, in terms of canonical isometric extensions of the channel family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$, and we provide an alternate proof based on basic properties of isometries and the spectral norm.
Theorem 23.
For a smooth family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of quantum channels, the following inequality holds for all $n\in\mathbb{N}$:
$\frac{1}{4}\sup_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}I_{F}\!\left(\theta;\left(\mathcal{P}_{\theta}^{(n)}\right)_{\theta\in\mathbb{R}}\right)$ $\leq n\inf_{\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{\left\|M_{H_{\theta}}\right\|^{2}+\left(n-1\right)\left\|V_{\theta}^{{\dagger}}M_{H_{\theta}}\right\|\left\|M_{H_{\theta}}\right\|\right\},$ (7.17)
where $V_{\theta}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ is a canonical isometric extension of $\mathcal{N}_{\theta}$ of the form in (5.73), the optimization is over every smooth family $\left(H_{\theta}\right)_{\theta\in\Theta}$ of Hermitian operators, and
$M_{H_{\theta}}\equiv\left(\partial_{\theta}V_{\theta}\right)-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta}.$ (7.18)
Proof.
Let $\hat{V}_{\theta}$ be defined as in (7.3). We can take isometric extensions of every step of the protocol $\mathcal{P}_{\theta}^{(n)}$ in (7.15) to write an isometric extension of $\mathcal{P}_{\theta}^{(n)}$ as follows:
$\left(I\otimes\hat{V}_{\theta}\right)A_{n-1}\cdots A_{2}\left(I\otimes\hat{V}_{\theta}\right)A_{1}\left(I\otimes\hat{V}_{\theta}\right),$ (7.19)
where $A_{i}$ is an isometric extension of $\mathcal{A}_{i}$. Suppressing implicit tensor products with identity in what follows, we apply Theorem 13 to conclude that
$\frac{1}{4}I_{F}\!\left(\theta;\left(\mathcal{P}_{\theta}^{(n)}\right)_{\theta\in\mathbb{R}}\right)$
$\leq\left\|\partial_{\theta}\left[\hat{V}_{\theta}A_{n-1}\cdots A_{2}\hat{V}_{\theta}A_{1}\hat{V}_{\theta}\right]\right\|^{2}$ (7.20)
$=\left\|\sum_{i=1}^{n}\hat{V}_{\theta}A_{n-1}\cdots A_{i+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right\|^{2}$ (7.21)
$=\left\|\begin{array}[]{c}\left(\sum_{j=1}^{n}\hat{V}_{\theta}A_{n-1}\cdots A_{j+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{j-1}\cdots A_{1}\hat{V}_{\theta}\right)^{{\dagger}}\times\\ \left(\sum_{i=1}^{n}\hat{V}_{\theta}A_{n-1}\cdots A_{i+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right)\end{array}\right\|$ (7.24)
$=\left\|\begin{array}[]{c}\left(\sum_{j=1}^{n}\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{\theta}^{{\dagger}}\right)\times\\ \left(\sum_{i=1}^{n}\hat{V}_{\theta}A_{n-1}\cdots A_{i+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right)\end{array}\right\|$ (7.27)
$=\left\|\sum_{i,j=1}^{n}\left(\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{\theta}^{{\dagger}}\right)\left(\hat{V}_{\theta}A_{n-1}\cdots A_{i+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right)\right\|$ (7.28)
$\leq\sum_{i=1}^{n}\left\|\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{i-1}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)A_{i+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{\theta}^{{\dagger}}\hat{V}_{\theta}A_{n-1}\cdots A_{i+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right\|$
$\ +\sum_{\begin{subarray}{c}i,j=1:\\ i\neq j\end{subarray}}^{n}\left\|\left(\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{\theta}^{{\dagger}}\right)\left(\hat{V}_{\theta}A_{n-1}\cdots A_{i+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right)\right\|.$ (7.29)
Consider that
$\sum_{i=1}^{n}\left\|\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{i-1}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)A_{i+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{\theta}^{{\dagger}}\hat{V}_{\theta}A_{n-1}\cdots A_{i+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right\|$
$=\sum_{i=1}^{n}\left\|\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{i-1}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right\|$ (7.30)
$\leq\sum_{i=1}^{n}\left\|\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{i-1}^{{\dagger}}A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right\|\left\|\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)\left(\partial_{\theta}\hat{V}_{\theta}\right)\right\|$ (7.31)
$=\sum_{i=1}^{n}\left\|\partial_{\theta}\hat{V}_{\theta}\right\|^{2}$ (7.32)
$=n\left\|\partial_{\theta}\hat{V}_{\theta}\right\|^{2}.$ (7.33)
For the cross terms where $i\neq j$, a generic term such that $i>j$ (without loss of generality) can be bounded from above as follows:
$\left\|\left(\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{\theta}^{{\dagger}}\right)\left(\hat{V}_{\theta}A_{n-1}\cdots A_{i+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right)\right\|$
$=\left\|\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)\hat{V}_{\theta}A_{j-1}\cdots A_{i+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right\|$ (7.34)
$\leq\left\|\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\right\|\left\|\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)\hat{V}_{\theta}\right\|\left\|A_{j-1}\cdots A_{i+1}\right\|\left\|\partial_{\theta}\hat{V}_{\theta}\right\|\left\|A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right\|$ (7.35)
$=\left\|\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)\hat{V}_{\theta}\right\|\left\|\partial_{\theta}\hat{V}_{\theta}\right\|$ (7.36)
$=\left\|\hat{V}_{\theta}^{{\dagger}}\partial_{\theta}\hat{V}_{\theta}\right\|\left\|\partial_{\theta}\hat{V}_{\theta}\right\|,$ (7.37)
where the inequality follows from submultiplicativity of the spectral norm. Thus, we conclude that
$\sum_{\begin{subarray}{c}i,j=1:\\ i\neq j\end{subarray}}^{n}\left\|\left(\hat{V}_{\theta}^{{\dagger}}A_{1}^{{\dagger}}\cdots A_{j-1}^{{\dagger}}\left(\partial_{\theta}\hat{V}_{\theta}^{{\dagger}}\right)A_{j+1}^{{\dagger}}\cdots A_{n-1}^{{\dagger}}\hat{V}_{\theta}^{{\dagger}}\right)\left(\hat{V}_{\theta}A_{n-1}\cdots A_{i+1}\left(\partial_{\theta}\hat{V}_{\theta}\right)A_{i-1}\cdots A_{1}\hat{V}_{\theta}\right)\right\|\\ \leq n\left(n-1\right)\left\|\hat{V}_{\theta}^{{\dagger}}\partial_{\theta}\hat{V}_{\theta}\right\|\left\|\partial_{\theta}\hat{V}_{\theta}\right\|.$ (7.38)
Applying (7.4) and Remark 14 and minimizing over every smooth family $\left(H_{\theta}\right)_{\theta\in\Theta}$ of Hermitian operators, we conclude the proof. ∎
Remark 24.
The upper bound in Theorem 22 does not exceed the upper bound in Theorem 23, consistent with the intuition that the SLD Fisher information for parallel protocols should not exceed the SLD Fisher information of adaptive protocols, given that parallel protocols are contained with the set of all adaptive protocols. This can be seen by comparing the objective functions in Theorems 22 and 23 and noting that the following inequality holds for every Hermitian operator $H_{\theta}$ and isometry $V_{\theta}$:
$\left\|V_{\theta}^{{\dagger}}M_{H_{\theta}}\right\|$ $=\left\|V_{\theta}V_{\theta}^{{\dagger}}M_{H_{\theta}}\right\|$ (7.39)
$\leq\left\|M_{H_{\theta}}\right\|$ (7.40)
where the equality follows because the spectral norm is invariant under left multiplication by the isometry $V_{\theta}$ and the inequality follows because $V_{\theta}V_{\theta}^{{\dagger}}$ is a projection and the spectral norm does not increase under the action of a projection.
8 Applications to query complexity
In this section, we present lower bounds on the error probabilities and query complexities of channel discrimination and estimation. With the main framework in place now, our proofs of these lower bounds follow somewhat directly from definitions and the upper bounds presented in Sections 6 and 7.
8.1 Lower bounds on query complexities of channel discrimination
We begin by establishing bounds on the performance of channel discrimination in the parallel and adaptive settings, in terms of the metric defined in (3.8):
Corollary 25.
Let $p\in\left(0,1\right)$ and set q=1pq=1-p. For channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ with canonical isometric extensions $V_{1}$ and $V_{2}$, the following bounds hold for the error probabilities of channel discrimination:
$1-\frac{p_{e,\|}\left(1-p_{e,\|}\right)}{pq}$ $\leq n\inf_{\begin{subarray}{c}W\in\mathbb{B}\end{subarray}}\left\{a_{W}+\left(n-1\right)b_{W}\right\},$ (8.1)
$1-\frac{p_{e,\operatorname{ad}}\left(1-p_{e,\operatorname{ad}}\right)}{pq}$ $\leq n\inf_{\begin{subarray}{c}W\in\mathbb{B}\end{subarray}}\left\{a_{W}+\left(n-1\right)\sqrt{a_{W}b_{W}}\right\},$ (8.2)
where
$a_{W}$ $\equiv 2\left\|I-\Re\!\left[M_{W}\right]\right\|,$ (8.3)
$b_{W}$ $\equiv\left\|I-M_{W}\right\|^{2},$ (8.4)
$p_{e,\|}$ $\equiv p_{e,\|}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n),$ (8.5)
$p_{e,\operatorname{ad}}$ $\equiv p_{e,\operatorname{ad}}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n),$ (8.6)
$M_{W}$ $\equiv V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2},$ (8.7)
and $p_{e,\|}$ and $p_{e,\operatorname{ad}}$ are defined in (3.1) and (3.4), respectively.
Proof.
Let us rewrite (3.1) and (3.4) as follows:
$\left\|p\mathcal{N}_{1}^{\otimes n}-q\mathcal{N}_{2}^{\otimes n}\right\|_{\diamond}$ $=1-2p_{e,\|},$ (8.8)
$\sup_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}\left\|p\mathcal{P}_{1}^{(n)}-q\mathcal{P}_{2}^{(n)}\right\|_{\diamond}$ $=1-2p_{e,\operatorname{ad}}.$ (8.9)
Now recall the following inequality holding for states $\rho$ and $\sigma$ (see, e.g., [CDL+25, Eqs. (F8)–(F13)]):
$\left\|p\rho-q\sigma\right\|_{1}^{2}\leq 1-4pq+4pqd_{B}^{2}(\rho,\sigma).$ (8.10)
By applying this to the output state of a parallel discrimination protocol and optimizing over input states, this then generalizes to
$\left\|p\mathcal{N}_{1}^{\otimes n}-q\mathcal{N}_{2}^{\otimes n}\right\|_{\diamond}^{2}\leq 1-4pq+4pqd_{B}^{2}(\mathcal{N}_{1}^{\otimes n},\mathcal{N}_{2}^{\otimes n}),$ (8.11)
which can be rewritten as
$1-\frac{p_{e,\|}\left(1-p_{e,\|}\right)}{pq}\leq d_{B}^{2}(\mathcal{N}_{1}^{\otimes n},\mathcal{N}_{2}^{\otimes n}),$ (8.12)
to which we apply Theorem 17 and conclude the proof of (8.1).
Similarly, for adaptive protocols, by applying (8.10) to the final output state of an adaptive channel discrimination protocol and optimizing over all such protocols, (8.10) generalizes to
$\left(\sup_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}\left\|p\mathcal{P}_{1}^{(n)}-q\mathcal{P}_{2}^{(n)}\right\|_{\diamond}\right)^{2}\leq 1-4pq+4pq\sup_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}d_{B}^{2}\!\left(\mathcal{P}_{1}^{(n)},\mathcal{P}_{2}^{(n)}\right),$ (8.13)
which can be rewritten as
$1-\frac{p_{e,\operatorname{ad}}\left(1-p_{e,\operatorname{ad}}\right)}{pq}\leq\sup_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}d_{B}^{2}\!\left(\mathcal{P}_{1}^{(n)},\mathcal{P}_{2}^{(n)}\right),$ (8.14)
to which we apply Theorem 19 and conclude the proof of (8.2). ∎
Corollary 26.
Let $p\in\left(0,1\right)$, set q=1pq=1-p, let $\varepsilon\in\left(0,\min\!\left\{p,q\right\}\right]$, and let $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ be quantum channels with canonical isometric extensions $V_{1}$ and $V_{2}$. Suppose that the conditions in Remark 2 do not hold. Then the following lower bounds hold for the query complexities of channel discrimination:
$n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\geq\left\lceil\sup_{W\in\mathbb{B}}\frac{b_{W}-a_{W}+\sqrt{\left(b_{W}-a_{W}\right)^{2}+4b_{W}\left(1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\right)}}{2b_{W}}\right\rceil$ (8.15)
$\geq\left\lceil\frac{1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}}{\inf_{W\in\mathbb{B}}\left\{a_{W}:M_{W}=I\right\}}\right\rceil,$ (8.16)
$n_{\operatorname{ad}}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\geq\left\lceil\sup_{\begin{subarray}{c}W\in\mathbb{B}\end{subarray}}\frac{c_{W}+\sqrt{c_{W}^{2}+4\sqrt{a_{W}b_{W}}\left(1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\right)}}{2\sqrt{a_{W}b_{W}}}\right\rceil$ (8.17)
$\geq\left\lceil\frac{1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}}{\inf_{W\in\mathbb{B}}\left\{a_{W}:M_{W}=I\right\}}\right\rceil,$ (8.18)
where
$a_{W}$ $\equiv 2\left\|I-\Re\!\left[M_{W}\right]\right\|,$ (8.19)
$b_{W}$ $\equiv\left\|I-M_{W}\right\|^{2},$ (8.20)
$c_{W}$ $\equiv\sqrt{a_{W}b_{W}}-a_{W},$ (8.21)
$M_{W}$ $\equiv V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}.$ (8.22)
Proof.
This is a direct consequence of Definition 1, Corollary 25 and Lemma 28 below, while noting that the function $x\mapsto x\left(1-x\right)$ is monotone increasing on the interval $\left[0,\frac{1}{2}\right]$. In more detail, we prove (8.15) and (8.16) and omit the proofs of (8.17) and (8.18), as they follow similarly.
Applying Definition 1, let $n\in\mathbb{N}$ be such that $p_{e,\|}\leq\varepsilon$, where we have adopted the shorthand $p_{e,\|}$, as defined in (8.5). Let $W\in\mathbb{B}$ be a contraction (i.e., $\left\|W\right\|\leq 1$). Then from the monotonicity of $x\mapsto x\left(1-x\right)$ on $\left[0,\frac{1}{2}\right]$ and Corollary 25, we conclude that
$1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}$ $\leq 1-\frac{p_{e,\|}\left(1-p_{e,\|}\right)}{pq},$ (8.23)
$\leq n\left(a_{W}+\left(n-1\right)b_{W}\right),$ (8.24)
where $a_{W}$ and $b_{W}$ are defined in (8.19) and (8.20), respectively. Applying Lemma 28 below, we conclude that
$n\geq\left\lceil\frac{b_{W}-a_{W}+\sqrt{\left(b_{W}-a_{W}\right)^{2}+4b_{W}\left(1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\right)}}{2b_{W}}\right\rceil.$ (8.25)
Since this lower bound holds for every contraction $W\in\mathbb{B}$, we can take a supremum over all such contractions to conclude that
$n\geq\sup_{\begin{subarray}{c}W\in\mathbb{B}\end{subarray}}\left\lceil\frac{b_{W}-a_{W}+\sqrt{\left(b_{W}-a_{W}\right)^{2}+4b_{W}\left(1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\right)}}{2b_{W}}\right\rceil.$ (8.26)
We then conclude (8.15) because this lower bound applies to all $n\in\mathbb{N}$ such that $p_{e,\|}\leq\varepsilon$.
In the case that there exists a contraction $W$ such that $M_{W}=I$, it follows that $b_{W}=0$, and (8.24) reduces to $1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\leq na_{W}$. Now applying Lemma 28 below, we conclude that
$n\geq\left\lceil\frac{1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}}{a_{W}}\right\rceil.$ (8.27)
Since this lower bound holds for every contraction $W\in\mathbb{B}$ satisfying $M_{W}=I$, we can take a supremum over all such contractions (which is equivalent to an infimum over the denominator) to conclude that
$n\geq\left\lceil\frac{1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}}{\inf_{W\in\mathbb{B}}\left\{a_{W}:M_{W}=I\right\}}\right\rceil.$ (8.28)
We then conclude (8.16) because this lower bound applies to all $n\in\mathbb{N}$ such that $p_{e,\|}\leq\varepsilon$. ∎
Remark 27.
The lower bounds in (8.16) and (8.18) are only applicable in the case that there exists a contraction $W$ satisfying $M_{W}=I$. If this is not the case, then
$\inf_{W\in\mathbb{B}}\left\{a_{W}:M_{W}=I\right\}=+\infty,$ (8.29)
implying that these lower bounds trivially evaluate to 0. Otherwise, the lower bounds in (8.16) and (8.18) do not trivially evaluate to zero, and the denominator can be efficiently computed by semi-definite optimization as follows:
$\inf_{W\in\mathbb{B}}\left\{a_{W}:M_{W}=I\right\}=\\ \inf_{\lambda\geq 0,W\in\mathbb{L}}\left\{\lambda:\lambda I\geq 2\left(I-\Re\!\left[M_{W}\right]\right)\geq-\lambda I,\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0,\ M_{W}=I\right\},$ (8.30)
where $M_{W}\equiv V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}$ and we appealed to (9.16) for $2\left\|I-\Re\!\left[M\right]\right\|$ and (9.8) for $\left\|W\right\|\leq 1$.
Lemma 28.
Suppose that the following inequality holds:
$n\left(a+\left(n-1\right)b\right)\geq c,$ (8.31)
where $a\geq b\geq 0$, $c\geq 0$, and $n\in\mathbb{N}$. Then the following lower bound holds on $n$:
$n\geq\begin{cases}\left\lceil\frac{b-a+\sqrt{\left(b-a\right)^{2}+4bc}}{2b}\right\rceil&:b>0\\ \left\lceil\frac{c}{a}\right\rceil&:b=0\end{cases}.$ (8.32)
Alternatively, suppose that the following inequality holds:
$n\left(a+\left(n-1\right)\sqrt{ab}\right)\geq c.$ (8.33)
Then the following lower bound holds on $n$:
$n\geq\begin{cases}\left\lceil\frac{\sqrt{ab}-a+\sqrt{\left(\sqrt{ab}-a\right)^{2}+4c\sqrt{ab}}}{2\sqrt{ab}}\right\rceil&:b>0\\ \left\lceil\frac{c}{a}\right\rceil&:b=0\end{cases}.$ (8.34)
Proof.
Consider that the inequality can be rewritten as
$bn^{2}+\left(a-b\right)n-c\geq 0$ (8.35)
Under the assumption that $b>0$, the parabola opens upwards, so that the inequality holds for $n\geq n_{+}$, where $n_{+}$ is the larger root of the quadratic equation $bn^{2}+\left(a-b\right)n-c=0$. By the quadratic formula, the two roots are given by
$n_{\pm}=\frac{-\left(a-b\right)\pm\sqrt{\left(a-b\right)^{2}+4bc}}{2b},$ (8.36)
and since $b>0$, the larger root is given by $n_{+}$. Thus, for $n\in\mathbb{N}$, the first lower bound in (8.32) holds. In the case that $b=0$, the inequality in (8.31) simplifies to $an\geq c$, from which we conclude the second expression in (8.32). The proof of (8.34) follows from the substitution $b\to\sqrt{ab}$. ∎
8.1.1 Finding lower bounds on query complexities of channel discrimination by binary search
In practice, the lower bounds in (8.15) and (8.17) of Corollary 26 may be difficult to compute, given that they are not obviously representable as convex optimization problems. However, it is possible to use the bounds from Corollary 25 in order to devise efficient algorithms for computing lower bounds on the query complexities of channel discrimination.
Let us begin with the parallel setting. For fixed $p\in\left(0,1\right)$, q=1pq=1-p, and $\varepsilon\in\left[0,\min\!\left\{p,q\right\}\right]$, if there exists an $n\in\mathbb{N}$ such that $p_{e,\|}\leq\varepsilon$, then, according to (8.24), the following inequality holds:
$1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\leq ng(n),$ (8.37)
where
$g(n)\coloneqq\inf_{\begin{subarray}{c}W\in\mathbb{B}\end{subarray}}\left\{a_{W}+\left(n-1\right)b_{W}\right\}.$ (8.38)
The function $g(n)$ can be calculated efficiently by means of a semi-definite program (see Proposition 39). Furthermore, $g(n)$ is a monotone non-decreasing function on $\mathbb{N}$ because, for every fixed contraction $W$, the following inequality holds:
$a_{W}+\left(n-1\right)b_{W}\leq a_{W}+nb_{W},$ (8.39)
given that $a_{W},b_{W}\geq 0$ for all $W$. Applying an infimum over every contraction $W$ to (8.39) then implies that $g(n)$ is monotone non-decreasing on $\mathbb{N}$, as claimed. We furthermore conclude that the function $n\mapsto ng(n)$ is monotone non-decreasing on $\mathbb{N}$.
Given this observation, we can find the smallest integer $n$ satisfying (8.37) (i.e., the query complexity $n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$) by binary search. Before doing so, it is helpful to have an upper bound on the query complexity. Toward this end, let us recall [NW25, Theorem 8], which implies the following upper bound on the query complexity $n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ by picking $s=\frac{1}{2}$ therein:
$n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\leq\left\lceil\frac{\ln\!\left(\frac{\sqrt{pq}}{\varepsilon}\right)}{Q_{\frac{1}{2}}(\mathcal{N}_{1}\|\mathcal{N}_{2})}\right\rceil,$ (8.40)
$\leq n_{\max}\coloneqq\left\lceil\frac{\ln\!\left(\frac{\sqrt{pq}}{\varepsilon}\right)}{-\ln\sqrt{F}(\mathcal{N}_{1},\mathcal{N}_{2})}\right\rceil$ (8.41)
where
$Q_{\frac{1}{2}}(\mathcal{N}_{1}\|\mathcal{N}_{2})$ $\coloneqq-\inf_{\psi_{RA}\in\mathbb{P}}\ln\operatorname{Tr}\!\left[\left[\left(\operatorname{id}\otimes\mathcal{N}_{1}\right)\left(\psi_{RA}\right)\right]^{\frac{1}{2}}\left[\left(\operatorname{id}\otimes\mathcal{N}_{2}\right)\left(\psi_{RA}\right)\right]^{\frac{1}{2}}\right],$ (8.42)
$F(\mathcal{N}_{1},\mathcal{N}_{2})$ $\coloneqq\inf_{\psi_{RA}\in\mathbb{P}}F\!\left(\left(\operatorname{id}\otimes\mathcal{N}_{1}\right)\left(\psi_{RA}\right),\left(\operatorname{id}\otimes\mathcal{N}_{2}\right)\left(\psi_{RA}\right)\right).$ (8.43)
The second inequality follows from reasoning similar to that used to arrive at [CDL+25, Corollary 8] (i.e., known inequalities between Holevo and Uhlmann fidelities, as recalled in [CDL+25, Eq. (A7)]). Furthermore, it is known how to compute the root fidelity of channels, $\sqrt{F}(\mathcal{N}_{1},\mathcal{N}_{2})$, efficiently by means of semi-definite optimization (see [KW21b, Proposition 55] or Proposition 37 below). As such, the upper bound $n_{\max}$ can be computed efficiently.
With this background in place, we now detail the following algorithm:
Algorithm 29.
The binary search algorithm for finding a lower bound on $n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ proceeds according to the following steps:
  1. 1.
    Initialize $n_{\ell}\leftarrow 1$, $n_{h}\leftarrow n_{\max}$.
  2. 2.
    While $n_{\ell}<n_{h}$,
    1. (a)
      Set $n_{m}\leftarrow\left\lfloor\frac{n_{\ell}+n_{h}}{2}\right\rfloor$.
    2. (b)
      Solve $g(n_{m})=\inf_{W\in\mathbb{B}}\left\{a_{W}+\left(n_{m}-1\right)b_{W}\right\}$ by semi-definite optimization (see Proposition 39).
    3. (c)
      If $n_{m}g(n_{m})\geq 1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}$, set $n_{h}\leftarrow n_{m}$. Otherwise, set $n_{\ell}\leftarrow n_{m}+1$.
  3. 3.
    Return $n_{\ell}$.
The total number of iterations for Algorithm 29 is equal to $O(\log n_{\max})$.
A similar procedure can be conducted for finding a lower bound on query complexity in the adaptive setting. Although the upper bound in (8.2) does not correspond to an SDP due to the presence of the square root, in practice, it can still be calculated efficiently by using the procedure outlined in Remark 41, which involves a combination of semi-definite optimization and one-dimensional grid search. For completeness, we provide the following algorithm for efficiently computing a lower bound on the query complexity of adaptive channel discrimination:
Algorithm 30.
The binary search algorithm for finding a lower bound on $n_{\operatorname{ad}}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ proceeds according to the following steps:
  1. 1.
    Initialize $n_{\ell}\leftarrow 1$, $n_{h}\leftarrow n_{\max}$.
  2. 2.
    While $n_{\ell}<n_{h}$,
    1. (a)
      Set $n_{m}\leftarrow\left\lfloor\frac{n_{\ell}+n_{h}}{2}\right\rfloor$.
    2. (b)
      Solve $g_{\operatorname{ad}}(n_{m})=\inf_{W\in\mathbb{B}}\left\{a_{W}+\left(n-1\right)\sqrt{a_{W}b_{W}}\right\}$ by the procedure outlined in Remark 41.
    3. (c)
      If $n_{m}g_{\operatorname{ad}}(n_{m})\geq 1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}$, set $n_{h}\leftarrow n_{m}$. Otherwise, set $n_{\ell}\leftarrow n_{m}+1$.
  3. 3.
    Return $n_{\ell}$.
Similar to Algorithm 29, the total number of iterations of Algorithm 30 is equal to $O(\log n_{\max})$.
Let us finally note that both of these algorithms are applicable and useful to determine lower bounds on the minimum number of queries needed to achieve perfect channel discrimination (i.e., $\varepsilon=0$). Indeed, Ref. [DFY09] established conditions under which two channels can be discriminated perfectly. By setting $\varepsilon=0$ in Algorithms 29 and 30, we can thus determine lower bounds on the query complexities of perfect channel discrimination in both the parallel and adaptive access models, respectively.
8.2 Lower bounds on query complexities of channel estimation
In this section, we leverage the lower bounds from (3.30)–(3.31), along with our lower bounds on the error probabilities and query complexities of channel discrimination from Section 8.1, in order to establish lower bounds on the minimax error probabilities and query complexities of channel estimation. We then establish asymptotic lower bounds on these quantities in terms of SLD Fisher information, arriving at results that generalize those from [MKSF+25, Eq. (53)] (see (8.61)–(8.64)). These asymptotic lower bounds bear similarity to quantum Cramer–Rao bounds for channel estimation, which have appeared in the context of quantum estimation theory (see, e.g., [DDKG12, KDD13, DDM14, ZJ21, KGADD23]).
Corollary 31.
For a parameterized family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of channels, the following bounds hold for the minimax error probability of channel estimation:
$1-4p_{e,\|}\left(1-p_{e,\|}\right)$ $\leq n\inf_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta,W\in\mathbb{B}\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}\left\{a_{W}^{\theta,\theta^{\prime}}+\left(n-1\right)b_{W}^{\theta,\theta^{\prime}}\right\},$ (8.44)
$1-4p_{e,\operatorname{ad}}\left(1-p_{e,\operatorname{ad}}\right)$ $\leq n\inf_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta,W\in\mathbb{B}\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}\left\{a_{W}^{\theta,\theta^{\prime}}+\left(n-1\right)\sqrt{a_{W}^{\theta,\theta^{\prime}}b_{W}^{\theta,\theta^{\prime}}}\right\},$ (8.45)
where
$p_{e,\|}$ $\equiv p_{e,\|}(\delta,\mathcal{F},n),$ (8.46)
$p_{e,\operatorname{ad}}$ $\equiv p_{e,\operatorname{ad}}(\delta,\mathcal{F},n),$ (8.47)
$a_{W}^{\theta,\theta^{\prime}}$ $\equiv 2\left\|I-\Re\!\left[M_{W}^{\theta,\theta^{\prime}}\right]\right\|,$ (8.48)
$b_{W}^{\theta,\theta^{\prime}}$ $\equiv\left\|I-M_{W}^{\theta,\theta^{\prime}}\right\|^{2},$ (8.49)
$M_{W}^{\theta,\theta^{\prime}}$ $\equiv V_{\theta}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{\theta^{\prime}},$ (8.50)
$V_{\theta}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ is a canonical isometric extension of $\mathcal{N}_{\theta}$ of the form in (5.73), and $p_{e,\|}$ and $p_{e,\operatorname{ad}}$ are defined in Definitions 3 and 5, respectively.
Proof.
These inequalities follow by combining (3.30) and (3.31) with Corollary 25, while noting that the function $x\mapsto x\left(1-x\right)$ is monotone increasing on the interval $\left[0,\frac{1}{2}\right]$. ∎
Corollary 32.
For a parameterized family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of channels and for all $\varepsilon\in\left(0,\frac{1}{2}\right]$, the following lower bounds hold for the query complexities of channel estimation:
$n_{\|}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq\left\lceil\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta,W\in\mathbb{B},\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}\frac{b_{W}^{\theta,\theta^{\prime}}-a_{W}^{\theta,\theta^{\prime}}+\sqrt{\left(b_{W}^{\theta,\theta^{\prime}}-a_{W}^{\theta,\theta^{\prime}}\right)^{2}+4b_{W}^{\theta,\theta^{\prime}}\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}}{2b_{W}^{\theta,\theta^{\prime}}}\right\rceil$ (8.51)
$\geq\left\lceil\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{W\in\mathbb{B}}\left\{a_{W}^{\theta,\theta^{\prime}}:M_{W}^{\theta,\theta^{\prime}}=I,\ \left|\theta-\theta^{\prime}\right|>2\delta,\ \theta,\theta^{\prime}\in\Theta\right\}}\right\rceil,$ (8.52)
$n_{\operatorname{ad}}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq\left\lceil\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta,W\in\mathbb{B},\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}\frac{c_{W}^{\theta,\theta^{\prime}}+\sqrt{\left(c_{W}^{\theta,\theta^{\prime}}\right)^{2}+4\sqrt{a_{W}^{\theta,\theta^{\prime}}b_{W}^{\theta,\theta^{\prime}}}\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}}{2\sqrt{a_{W}^{\theta,\theta^{\prime}}b_{W}^{\theta,\theta^{\prime}}}}\right\rceil$ (8.53)
$\geq\left\lceil\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{W\in\mathbb{B}}\left\{a_{W}^{\theta,\theta^{\prime}}:M_{W}^{\theta,\theta^{\prime}}=I,\ \left|\theta-\theta^{\prime}\right|>2\delta,\ \theta,\theta^{\prime}\in\Theta\right\}}\right\rceil,$ (8.54)
where
$a_{W}^{\theta,\theta^{\prime}}$ $\equiv 2\left\|I-\Re\!\left[M_{W}^{\theta,\theta^{\prime}}\right]\right\|,$ (8.55)
$b_{W}^{\theta,\theta^{\prime}}$ $\equiv\left\|I-M_{W}^{\theta,\theta^{\prime}}\right\|^{2},$ (8.56)
$c_{W}^{\theta,\theta^{\prime}}$ $\equiv\sqrt{a_{W}^{\theta,\theta^{\prime}}b_{W}^{\theta,\theta^{\prime}}}-a_{W}^{\theta,\theta^{\prime}},$ (8.57)
$M_{W}^{\theta,\theta^{\prime}}$ $\equiv V_{\theta}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{\theta^{\prime}},$ (8.58)
and $V_{\theta}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ is a canonical isometric extension of $\mathcal{N}_{\theta}$ of the form in (5.73).
Proof.
These inequalities follow by combining (3.41) and (3.42) with Corollary 26, while noting that the function $x\mapsto x\left(1-x\right)$ is monotone increasing on the interval $\left[0,\frac{1}{2}\right]$. ∎
Remark 33.
Similar to what was stated in Remark 27, the lower bounds in (8.52) and (8.54) are only applicable in the case that, for some $\theta,\theta^{\prime}\in\Theta$ satisfying $\left|\theta-\theta^{\prime}\right|>2\delta$, there exists $W\in\mathbb{B}$ such that $M_{W}^{\theta,\theta^{\prime}}=I$. Otherwise these lower bounds are trivially equal to 0.
Finally, we obtain asymptotic expressions for bounds on the minimax error probabilities and query complexities by appealing to (8.12), (8.14), (3.30), (3.31), and Theorems 16, 22, and 23.
Corollary 34.
For a smooth family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)$ of channels and $\varepsilon\in\left(0,\frac{1}{2}\right]$, the following bounds hold for the minimax error probabilities and query complexities of channel estimation:
$1-4p_{e,\|}\left(1-p_{e,\|}\right)$ $\leq\frac{\delta^{2}}{4}n\inf_{\begin{subarray}{c}\theta\in\Theta,\\ \left(H_{\theta}\right)_{\theta\in\Theta}\end{subarray}}\left\{a_{\theta}+\left(n-1\right)b_{\theta}\right\}+o(\delta^{2}),$ (8.59)
$1-4p_{e,\operatorname{ad}}\left(1-p_{e,\operatorname{ad}}\right)$ $\leq\frac{\delta^{2}}{4}n\inf_{\begin{subarray}{c}\theta\in\Theta,\\ \left(H_{\theta}\right)_{\theta\in\Theta}\end{subarray}}\left\{a_{\theta}+\left(n-1\right)\sqrt{a_{\theta}b_{\theta}}\right\}+o(\delta^{2}),$ (8.60)
$n_{\|}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{b_{\theta}\right\}}}\right)$ (8.61)
$\geq O\!\left(\frac{1}{\delta^{2}}\left(\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta}\left\{a_{\theta}:V_{\theta}^{{\dagger}}M_{H_{\theta}}=0,\ \left(H_{\theta}\right)_{\theta\in\Theta}\right\}}\right)\right),$ (8.62)
$n_{\operatorname{ad}}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{\sqrt{a_{\theta}b_{\theta}}\right\}}}\right)$ (8.63)
$\geq O\!\left(\frac{1}{\delta^{2}}\left(\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta}\left\{a_{\theta}:V_{\theta}^{{\dagger}}M_{H_{\theta}}=0,\ \left(H_{\theta}\right)_{\theta\in\Theta}\right\}}\right)\right),$ (8.64)
where
$p_{e,\|}$ $\equiv p_{e,\|}(\delta,\mathcal{F},n),$ (8.65)
$p_{e,\operatorname{ad}}$ $\equiv p_{e,\operatorname{ad}}(\delta,\mathcal{F},n),$ (8.66)
$a_{\theta}$ $\equiv\left\|M_{H_{\theta}}\right\|^{2},$ (8.67)
$b_{\theta}$ $\equiv\left\|V_{\theta}^{{\dagger}}M_{H_{\theta}}\right\|^{2},$ (8.68)
$M_{H_{\theta}}$ $\equiv\left(\partial_{\theta}V_{\theta}\right)-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta},$ (8.69)
$V_{\theta}\colon\mathbb{C}^{d_{A}}\mapsto\mathbb{C}^{d_{E}}\otimes\mathbb{C}^{d_{B}}$ is a canonical isometric extension of $\mathcal{N}_{\theta}$ of the form in (5.73) and the optimizations are over every smooth family $\left(H_{\theta}\right)_{\theta\in\Theta}$ of Hermitian operators.
Proof.
The inequalities in (8.59) and (8.60) are indeed direct consequences of (8.12), (8.14), (3.30), (3.31), and Theorem 16, as stated.
To obtain the query complexity lower bounds in (8.15)–(8.18), we apply Lemma 28. Let us begin with the parallel setting. Let $\theta\in\Theta$ and let $\left(H_{\theta}\right)_{\theta\in\Theta}$ be a smooth family of Hermitian operators. Then the following bound holds, as a consequence of (8.59):
$1-4p_{e,\|}\left(1-p_{e,\|}\right)\leq\frac{n}{4}\left(\delta^{2}a_{\theta}+\left(n-1\right)\delta^{2}b_{\theta}\right)+o(\delta^{2}).$ (8.70)
Ignoring the $o(\delta^{2})$ term, as we are considering the limit $\delta\to 0$, and applying Lemma 28, we conclude that
$n$ $\geq 4\left(\frac{\delta^{2}b_{\theta}-\delta^{2}a_{\theta}+\sqrt{\left(\delta^{2}b_{\theta}-\delta^{2}a_{\theta}\right)^{2}+4\delta^{2}b_{\theta}\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}}{2\delta^{2}b_{\theta}}\right)$ (8.71)
$=\frac{4}{\delta}\left(\frac{\delta b_{\theta}-\delta a_{\theta}+\sqrt{\left(\delta b_{\theta}-\delta a_{\theta}\right)^{2}+4b_{\theta}\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}}{2b_{\theta}}\right)$ (8.72)
$=O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{b_{\theta}}}\right).$ (8.73)
Since the bound holds for every $\theta\in\Theta$ and $\left(H_{\theta}\right)_{\theta\in\Theta}$ and every $n\in\mathbb{N}$ such that $p_{e,\|}\leq\varepsilon$, we can take an infimum over both to conclude that
$n_{\|}^{\star}(\varepsilon,\delta,\mathcal{F})\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{b_{\theta}\right\}}}\right),$ (8.74)
which proves (8.61).
In the case that there exists $\theta\in\Theta$ and a smooth family $\left(H_{\theta}\right)_{\theta\in\Theta}$ of Hermitian operators such that $b_{\theta}=0$, then the upper bound in (8.70) becomes
$1-4p_{e,\|}\left(1-p_{e,\|}\right)\leq\frac{n}{4}\delta^{2}a_{\theta}+o(\delta^{2}).$ (8.75)
Now flipping this inequality around and applying a similar line of reasoning, we conclude (8.62).
We now argue for the adaptive setting, with a similar line of reasoning. Let $\theta\in\Theta$ and let $\left(H_{\theta}\right)_{\theta\in\Theta}$ be a smooth family of Hermitian operators. Then the following bound holds, as a consequence of (8.60):
$1-4p_{e,\operatorname{ad}}\left(1-p_{e,\operatorname{ad}}\right)$ $\leq\frac{n}{4}\delta^{2}\left(a_{\theta}+\left(n-1\right)\sqrt{a_{\theta}b_{\theta}}\right)+o(\delta^{2})$ (8.76)
$=\frac{n}{4}\left(\delta^{2}a_{\theta}+\left(n-1\right)\sqrt{\left(\delta^{2}a_{\theta}\right)\left(\delta^{2}b_{\theta}\right)}\right)+o(\delta^{2})$ (8.77)
Ignoring the $o(\delta^{2})$ term, as we are considering the limit $\delta\to 0$, and applying Lemma 28, we conclude that
$n$ $\geq 4\left(\frac{\sqrt{\delta^{2}a_{\theta}\delta^{2}b_{\theta}}-\delta^{2}a_{\theta}+\sqrt{\left(\sqrt{\delta^{2}a_{\theta}\delta^{2}b_{\theta}}-\delta^{2}a_{\theta}\right)^{2}+4\sqrt{\delta^{2}a_{\theta}\delta^{2}b_{\theta}}\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}}{2\sqrt{\delta^{2}a_{\theta}\delta^{2}b_{\theta}}}\right)$ (8.78)
$=\frac{4}{\delta}\left(\frac{\sqrt{\delta a_{\theta}\delta b_{\theta}}-\delta a_{\theta}+\sqrt{\left(\sqrt{\delta a_{\theta}\delta b_{\theta}}-\delta a_{\theta}\right)^{2}+4\sqrt{a_{\theta}b_{\theta}}\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}}{2\sqrt{a_{\theta}b_{\theta}}}\right)$ (8.79)
$=O\!\left(\frac{1}{\delta}\sqrt{\frac{\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}{\sqrt{a_{\theta}b_{\theta}}}}\right).$ (8.80)
Since the bound holds for every $\theta\in\Theta$ and $\left(H_{\theta}\right)_{\theta\in\Theta}$ and every $n\in\mathbb{N}$ such that $p_{e,\operatorname{ad}}\leq\varepsilon$, we can take an infimum over both to conclude that
$n_{\operatorname{ad}}^{\star}(\varepsilon,\delta,\mathcal{F})\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{\sqrt{a_{\theta}b_{\theta}}\right\}}}\right),$ (8.81)
which is (8.63).
In the case that there exists $\theta\in\Theta$ and a smooth family $\left(H_{\theta}\right)_{\theta\in\Theta}$ of Hermitian operators such that $b_{\theta}=0$, then the upper bound in (8.77) becomes
$1-4p_{e,\|}\left(1-p_{e,\|}\right)\leq\frac{n}{4}\delta^{2}a_{\theta}+o(\delta^{2}).$ (8.82)
Now flipping this inequality around and applying a similar line of reasoning, we conclude (8.64). ∎
Remark 35.
The lower bounds in (8.62) and (8.64) are only applicable in the case that there exists $\theta\in\Theta$ and a Hermitian operator $H_{\theta}$ such that $V_{\theta}^{{\dagger}}M_{H_{\theta}}=0$. If this is the case, then Heisenberg scaling does not occur. If this is not the case, then
$\inf_{\theta\in\Theta}\left\{a_{\theta}:V_{\theta}^{{\dagger}}M_{H_{\theta}}=0,\ \left(H_{\theta}\right)_{\theta\in\Theta}\right\}=+\infty,$ (8.83)
implying that these lower bounds trivially evaluate to 0. Otherwise, the lower bounds in (8.62) and (8.64) do not trivially evaluate to zero, and the denominator can be efficiently computed by a combination of one-dimensional grid search over $\theta\in\Theta$ and semi-definite optimization as follows:
$\inf_{H_{\theta}\in\mathbb{H}}\left\{a_{\theta}:V_{\theta}^{{\dagger}}M_{H_{\theta}}=0\right\}=\\ \inf_{\lambda\geq 0}\left\{\lambda:\begin{bmatrix}\lambda I&M_{H_{\theta}}\\ M_{H_{\theta}}^{{\dagger}}&I\end{bmatrix}\geq 0,\ M_{H_{\theta}}=\left(\partial_{\theta}V_{\theta}\right)-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta},\ V_{\theta}^{{\dagger}}M_{H_{\theta}}=0\right\},$ (8.84)
where we appealed to (9.13) for $a_{\theta}=\left\|M_{H_{\theta}}\right\|^{2}$.
Remark 36.
Similar to the approaches outlined in Section 8.1.1, lower bounds on the query complexities of channel estimation can be efficiently computed by a combination of semi-definite optimization and grid search. Indeed, from the bounds in Corollary 31 and algorithms similar to Algorithms 29 and 30, combined with an outer search over $\theta,\theta^{\prime}\in\Theta$ satisfying $\left|\theta-\theta^{\prime}\right|>2\delta$, one can efficiently compute lower bounds on the query complexities of channel estimation. One can obtain looser bounds at lower computational cost by performing an outer search over just $\theta\in\Theta$ and setting $\theta^{\prime}=\theta+2\delta$. For sufficiently small $\delta$, one could alternatively obtain lower bounds on the query complexities of channel estimation using the bounds from Corollary 34 and algorithms similar to Algorithms 29 and 30, combined with an outer search over $\theta\in\Theta$.
9 Optimization of channel discrimination and estimation measures
In this section, we show how to compute all of the channel measures in our paper in terms of semi-definite optimization alone or semi-definite optimization combined with grid search. One of the main tools that we employ for this purpose is the Schur complement lemma.
9.1 Schur complement lemma and implications for optimization
The Schur complement lemma states that the following equivalence holds:
$\begin{bmatrix}K&M\\ M^{{\dagger}}&L\end{bmatrix}>0\qquad\Longleftrightarrow\qquad L>0,\quad K>M^{{\dagger}}L^{-1}M.$ (9.1)
It can be extended to positive semi-definite constraints in the following way:
$\begin{bmatrix}K&M\\ M^{{\dagger}}&L\end{bmatrix}\geq 0\qquad$ $\Longleftrightarrow\qquad\begin{bmatrix}K+\delta I&M\\ M^{{\dagger}}&L+\varepsilon I\end{bmatrix}>0\quad\forall\delta,\varepsilon>0$ (9.2)
$\Longleftrightarrow\qquad L+\varepsilon I>0,\quad K+\delta I>M^{{\dagger}}\left(L+\varepsilon I\right)^{-1}M\quad\forall\delta,\varepsilon>0$ (9.3)
$\Longleftrightarrow\qquad L\geq 0,\quad K\geq M^{{\dagger}}\left(L+\varepsilon I\right)^{-1}M\quad\forall\varepsilon>0.$ (9.4)
If $\operatorname{im}(M)\perp\ker(L)$, then
$\lim_{\varepsilon\to 0^{+}}M^{{\dagger}}\left(L+\varepsilon I\right)^{-1}M=M^{{\dagger}}L^{-1}M,$ (9.5)
where the inverse on the right-hand side is understood to be a generalized inverse, taken on the support of $L$, so that
$\begin{bmatrix}K&M\\ M^{{\dagger}}&L\end{bmatrix}\geq 0\qquad\Longleftrightarrow\qquad L\geq 0,\quad K\geq M^{{\dagger}}L^{-1}M,$ (9.6)
whenever $\operatorname{im}(M)\perp\ker(L)$.
The Schur complement lemma leads to the following conclusions for $A\in\mathbb{L}$:
$\begin{bmatrix}\lambda I&A\\ A^{{\dagger}}&I\end{bmatrix}$ $\geq 0\qquad\Longleftrightarrow\qquad\lambda I\geq A^{{\dagger}}A,$ (9.7)
$\begin{bmatrix}I&A\\ A^{{\dagger}}&I\end{bmatrix}$ $\geq 0\qquad\Longleftrightarrow\qquad 1\geq\left\|A\right\|,$ (9.8)
$\begin{bmatrix}\lambda I&A\\ A^{{\dagger}}&\lambda I\end{bmatrix}$ $\geq 0\qquad\Longleftrightarrow\qquad\lambda\geq 0,\quad\lambda^{2}I\geq A^{{\dagger}}A,$ (9.9)
with the latter condition following because
$\begin{bmatrix}\lambda I&A\\ A^{{\dagger}}&\lambda I\end{bmatrix}\geq 0\qquad\Longleftrightarrow$ $\qquad\lambda I\geq 0,\quad\lambda I\geq A^{{\dagger}}\left(\lambda I+\varepsilon I\right)^{-1}A\quad\forall\varepsilon>0,$ (9.10)
$\Longleftrightarrow$ $\qquad\lambda\geq 0,\quad\lambda(\lambda+\varepsilon)I\geq A^{{\dagger}}A\quad\forall\varepsilon>0,$ (9.11)
$\Longleftrightarrow$ $\qquad\lambda\geq 0,\quad\lambda^{2}I\geq A^{{\dagger}}A.$ (9.12)
Given that $\left\|A\right\|^{2}=\lambda_{\max}(A^{{\dagger}}A)$ and $\left\|A\right\|=\sqrt{\lambda_{\max}(A^{{\dagger}}A)}$, we arrive at the following semi-definite programs (SDPs):
$\left\|A\right\|^{2}$ $=\inf_{\lambda\geq 0}\left\{\lambda:\begin{bmatrix}\lambda I&A\\ A^{{\dagger}}&I\end{bmatrix}\geq 0\right\},$ (9.13)
$\left\|A\right\|$ $=\inf_{\lambda\geq 0}\left\{\lambda:\begin{bmatrix}\lambda I&A\\ A^{{\dagger}}&\lambda I\end{bmatrix}\geq 0\right\}.$ (9.14)
If the matrix of interest, now denoted by $B$, is either positive semi-definite or Hermitian, SDP characterizations of $\left\|B\right\|$ can be simplified along the following lines:
$B$ $\geq 0\quad\implies\quad\left\|B\right\|=\inf_{\lambda\geq 0}\left\{\lambda:B\leq\lambda I\right\},$ (9.15)
$B$ $\in\mathbb{H}\quad\implies\quad\left\|B\right\|=\inf_{\lambda\geq 0}\left\{\lambda:-\lambda I\leq B\leq\lambda I\right\}.$ (9.16)
9.2 Optimization of quantities based on Bures distance of channels
In this section, we delineate methods for optimizing various channel distinguishability measures related to the Bures distance.
SDPs for the root fidelity of channels were established by [YF17, Eq. (7)] and [KW21b, Proposition 55]. Proposition 37 gives an alternative semi-definite program for computing the root fidelity of channels.
Proposition 37.
The root fidelity of channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ can be expressed as the following semi-definite program:
$\sqrt{F}(\mathcal{N}_{1},\mathcal{N}_{2})=\sup_{\lambda\geq 0,W\in\mathbb{L}}\left\{\lambda:\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\geq\lambda I,\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0\right\}.$ (9.17)
Proof.
Beginning with the developments in (4.49) and (4.50)–(4.55), consider that
$\sqrt{F}(\mathcal{N}_{1},\mathcal{N}_{2})$ $=\sup_{W\in\mathbb{B}}\inf_{\rho}\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]$ (9.18)
$=\sup_{W\in\mathbb{B}}\inf_{\rho}\operatorname{Tr}\!\left[\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\rho\right]$ (9.19)
$=\sup_{W\in\mathbb{B}}\sup_{\lambda\in\mathbb{R}}\left\{\lambda:\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\geq\lambda I\right\}$ (9.20)
$=\sup_{W\in\mathbb{B}}\sup_{\lambda\geq 0}\left\{\lambda:\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\geq\lambda I\right\}$ (9.21)
$=\sup_{\lambda\geq 0,W\in\mathbb{L}}\left\{\lambda:\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\geq\lambda I,\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0\right\},$ (9.22)
where we employed (9.8). ∎
Proposition 38.
The squared Bures distance of channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ can be expressed as the following SDP:
$d_{B}^{2}(\mathcal{N}_{1},\mathcal{N}_{2})=\inf_{\begin{subarray}{c}\lambda\in\mathbb{R},W\in\mathbb{L}\end{subarray}}\left\{\begin{array}[]{c}\lambda:\lambda I\geq 2\left(I-\Re\!\left[M\right]\right)\geq-\lambda I,\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0,\\ M=V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\end{array}\right\}.$ (9.23)
Proof.
Starting from (4.41) and applying (9.16) for $2\left\|I-\Re\!\left[M\right]\right\|$ and (9.8) for $\left\|W\right\|\leq 1$, consider that
$d_{B}^{2}(\mathcal{N}_{1},\mathcal{N}_{2})$ $=\inf_{W\in\mathbb{B}}\left\|2\left(I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right)\right\|$ (9.24)
$=\inf_{\begin{subarray}{c}\lambda\in\mathbb{R},\\ W\in\mathbb{L}\end{subarray}}\left\{\begin{array}[]{c}\lambda:\lambda I\geq 2\left(I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right)\geq-\lambda I,\\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0\end{array}\right\},$ (9.27)
thus concluding the proof. ∎
Proposition 39.
The upper bound in Theorem 17 for parallel discrimination of channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ can be expressed as the following SDP:
$\inf_{W,M\in\mathbb{L}:\left\|W\right\|\leq 1}\left\{2\left\|I-\Re\!\left[M\right]\right\|+\left(n-1\right)\left\|I-M\right\|^{2}:M=V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right\}\\ =\inf_{\begin{subarray}{c}\lambda,\mu\geq 0,\\ W,M\in\mathbb{L}\end{subarray}}\left\{\begin{array}[]{c}\lambda+\left(n-1\right)\mu:\lambda I\geq 2\left(I-\Re\!\left[M\right]\right)\geq-\lambda I,\ \begin{bmatrix}\mu I&I-M\\ I-M^{{\dagger}}&I\end{bmatrix}\geq 0,\\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0,\ M=V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\end{array}\right\},$ (9.28)
Proof.
Applying (9.16) for $2\left\|I-\Re\!\left[M\right]\right\|$, (9.13) for $\left\|I-M\right\|^{2}$, and (9.8) for $\left\|W\right\|\leq 1$, consider that
$\inf_{W\in\mathbb{B}}\left\{\left\|2\left(I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right)\right\|+\left(n-1\right)\left\|I-V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right\|^{2}\right\}\\ =\inf_{\begin{subarray}{c}\lambda,\mu\geq 0,\\ W\in\mathbb{L}\end{subarray}}\left\{\begin{array}[]{c}\lambda+\left(n-1\right)\mu:\\ \lambda I\geq 2\left(I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right)\geq-\lambda I,\\ \begin{bmatrix}\mu I&I-V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\\ I-\left(V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right)^{{\dagger}}&I\end{bmatrix}\geq 0,\\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0\end{array}\right\},$ (9.29)
thus concluding the proof. ∎
Proposition 40.
The upper bound in Theorem 19 for adaptive discrimination of channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ can be expressed in the following way:
$\inf_{\begin{subarray}{c}W,M\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\left\{2\left\|I-\Re\!\left[M\right]\right\|+\left(n-1\right)\left(2\left\|I-\Re\!\left[M\right]\right\|\right)^{\frac{1}{2}}\left\|I-M\right\|:M=V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right\}\\ =\inf_{\nu\in\left(0,1\right]}\inf_{\begin{subarray}{c}\lambda,\mu\geq 0,\\ W\in\mathbb{L}\end{subarray}}\left\{\begin{array}[]{c}\left(1+\frac{n-1}{2}\nu\right)\lambda+\left(\frac{n-1}{2\nu}\right)\mu:\\ \lambda I\geq 2\left(I-\Re\!\left[M\right]\right)\geq-\lambda I,\ \begin{bmatrix}\mu I&I-M\\ I-M^{{\dagger}}&I\end{bmatrix}\geq 0,\\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0,\ M=V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2},\end{array}\right\}.$ (9.30)
Proof.
Consider that, for $a,b\geq 0$, such that $b\leq a$, the following equality holds:
$\sqrt{ab}=\frac{1}{2}\inf_{\nu\in\left(0,1\right]}\left\{\nu a+\frac{b}{\nu}\right\},$ (9.31)
while noting that the optimal value of $\nu$ is $\sqrt{\frac{b}{a}}$. Keeping the constraint $M=V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}$ implicit, applying (9.31), while noting Remark 21, it follows that
$\inf_{\begin{subarray}{c}W,M\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\left\{2\left\|I-\Re\!\left[M\right]\right\|+\left(n-1\right)\left(2\left\|I-\Re\!\left[M\right]\right\|\right)^{\frac{1}{2}}\left\|I-M\right\|\right\}$
$=\inf_{\begin{subarray}{c}W,M\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\left\{2\left\|I-\Re\!\left[M\right]\right\|+\left(n-1\right)\sqrt{2\left\|I-\Re\!\left[M\right]\right\|\left\|I-M\right\|^{2}}\right\}$ (9.32)
$=\inf_{\nu\in\left(0,1\right]}\inf_{\begin{subarray}{c}W,M\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\left\{2\left\|I-\Re\!\left[M\right]\right\|+\left(\frac{n-1}{2}\right)\left(\nu\left(2\left\|I-\Re\!\left[M\right]\right\|\right)+\frac{\left\|I-M\right\|^{2}}{\nu}\right)\right\}$ (9.33)
$=\inf_{\nu\in\left(0,1\right]}\inf_{\begin{subarray}{c}W,M\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\left\{1+\frac{n-1}{2}\nu\left(2\left\|I-\Re\!\left[M\right]\right\|\right)+\left(\frac{n-1}{2}\right)\frac{\left\|I-M\right\|^{2}}{\nu}\right\}.$ (9.34)
Applying (9.16) for $2\left\|I-\Re\!\left[M\right]\right\|$ and (9.13) for $\left\|I-M\right\|^{2}$, we conclude (9.30). ∎
Remark 41.
In practice, the optimization in (9.30) can be computed efficiently, even though it is not an SDP. Indeed, (9.30) can be rewritten in the following way:
$\inf_{\nu\in\left(0,1\right]}J(\nu),$ (9.35)
where
$J(\nu)\coloneqq\inf_{\begin{subarray}{c}\lambda,\mu\geq 0,\\ W\in\mathbb{L}\end{subarray}}\left\{\begin{array}[]{c}\left(1+\frac{n-1}{2}\nu\right)\lambda+\left(\frac{n-1}{2\nu}\right)\mu:\\ \lambda I\geq 2\left(I-\Re\!\left[M\right]\right)\geq-\lambda I,\ \begin{bmatrix}\mu I&I-M\\ I-M^{{\dagger}}&I\end{bmatrix}\geq 0,\\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0,\ M=V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2},\end{array}\right\}.$ (9.36)
For each fixed $\nu\in\left(0,1\right]$, the function $J(\nu)$ can be computed efficiently because it is an SDP. As such, one can perform a global one-dimensional line search for the value of $\nu\in\left(0,1\right]$ that minimizes $J(\nu)$. The global search is inexpensive in practice because $\nu$ is one-dimensional, and there are well known methods for performing this step [PTVF07]. In more detail, the outer one-dimensional optimization over $\nu$ is nonconvex in general. However, since the problem is scalar, a global optimum can be obtained efficiently by a coarse-to-fine search strategy: we can first sample $J(\nu)$ on a logarithmic grid, and then refine the best candidates using Brent’s derivative-free minimization algorithm [Bre73].
9.3 Optimization of quantities based on SLD Fisher information
In this section, we provide various methods for optimizing quantities involving the SLD Fisher information. We note that, recently, a numerical package is available for this purpose [DKDD25]. In some cases, the approaches we outline below differ from those given in prior works [KGADD23, DKDD25].
Let us begin with the SLD Fisher information of channels, while noting that a similar SDP for it was proposed in [DDKG12, Eqs. (21)–(22)].
Proposition 42.
The SLD Fisher information of the smooth channel family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ can be expressed as the following SDP:
$\frac{1}{4}I_{F}\!\left(\theta;\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}\right)$ $=\inf_{H_{\theta}\in\mathbb{H}}\left\|M_{H_{\theta}}\right\|^{2}$ (9.37)
$=\inf_{\lambda\geq 0,H_{\theta}\in\mathbb{H}}\left\{\lambda:\begin{bmatrix}\lambda I&M_{H_{\theta}}\\ M_{H_{\theta}}^{{\dagger}}&I\end{bmatrix}\geq 0\right\},$ (9.38)
where
$M_{H_{\theta}}\equiv\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta}.$ (9.39)
Proof.
This is a direct consequence of (5.76) and (9.13), while incorporating the optimization over $H_{\theta}\in\mathbb{H}$. ∎
The following SDP is similar to that given in [KDD13, Appendix E]:
Proposition 43.
For all $n\in\mathbb{N}$, the upper bound in Theorem 22 for parallel estimation of the smooth channel family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ can be expressed as the following SDP:
$\inf_{H_{\theta}\in\mathbb{H}}\left\{\left\|M_{H_{\theta}}\right\|^{2}+\left(n-1\right)\left\|V_{\theta}^{{\dagger}}M_{H_{\theta}}\right\|^{2}\right\}\\ =\inf_{\begin{subarray}{c}\lambda,\mu\geq 0,\\ H_{\theta}\in\mathbb{H}\end{subarray}}\left\{\begin{array}[]{c}\lambda+\left(n-1\right)\mu:\begin{bmatrix}\lambda I&M_{H_{\theta}}\\ M_{H_{\theta}}^{{\dagger}}&I\end{bmatrix}\geq 0,\ \begin{bmatrix}\mu I&V_{\theta}^{{\dagger}}M_{H_{\theta}}\\ M_{H_{\theta}}^{{\dagger}}V_{\theta}&I\end{bmatrix}\geq 0,\end{array}\right\},$ (9.40)
where
$M_{H_{\theta}}\equiv\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta}.$ (9.41)
Proof.
This is a direct consequence of (9.13), while incorporating the optimization over $H_{\theta}\in\mathbb{H}$. ∎
Proposition 44.
For all $n\in\mathbb{N}$, the upper bound in Theorem 23 for adaptive estimation of the smooth channel family $\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ can be expressed in the following way:
$\inf_{H_{\theta}\in\mathbb{H}}\left\{\left\|M_{H_{\theta}}\right\|^{2}+\left(n-1\right)\left\|M_{H_{\theta}}\right\|\left\|V_{\theta}^{{\dagger}}M_{H_{\theta}}\right\|\right\}=\\ \inf_{\nu\in\left(0,1\right]}\inf_{\begin{subarray}{c}\lambda,\mu\geq 0,\\ H_{\theta}\in\mathbb{H}\end{subarray}}\left\{\left(1+\frac{n-1}{2}\nu\right)\lambda+\left(\frac{n-1}{2\nu}\right)\mu:\begin{bmatrix}\lambda I&M_{H_{\theta}}\\ M_{H_{\theta}}&I\end{bmatrix}\geq 0,\begin{bmatrix}\mu I&V_{\theta}^{{\dagger}}M_{H_{\theta}}\\ M_{H_{\theta}}^{{\dagger}}V_{\theta}&I\end{bmatrix}\geq 0\right\},$ (9.42)
where
$M_{H_{\theta}}\equiv\partial_{\theta}V_{\theta}-i\left(H_{\theta}\otimes I_{B}\right)V_{\theta}.$ (9.43)
Proof.
The proof follows a similar line of reasoning as given for Proposition 40. ∎
By the same line of reasoning given in Remark 41, the optimization in (9.42) can be computed efficiently by a combination of semi-definite optimization and one-dimensional grid search. This approach differs from that outlined in [KGADD23, Appendix D].
10 Conclusion
In summary, we provided a unified treatment for quantum channel discrimination and estimation, and we established lower bounds on the error probabilities and query complexities for these tasks, in both the parallel and adaptive access models. As highlighted throughout, an advantage of our approach is that we formulate all statements and proofs in terms of isometric extensions of quantum channels, so that many proofs rely on basic properties of isometries and the spectral norm. Along the way, we also revisited many foundational statements in the quantum estimation theory literature, providing alternative proofs that may enhance general understanding of them.
Going forward, there are a number of questions that can be addressed using our framework. First, can one easily generalize the various bounds presented here to the energy-constrained case? There has already been some work on this topic [CY25], and we suspect that the operator $E$-norm of [Shi20], combined with our approach here, could be useful for this purpose. Additionally, we wonder about extending the results to obtain lower bounds on the query complexity of multiple channel discrimination. Bounds for this task are available in [NW25, Theorem 19], but the approach of our paper, possibly combined with the notion of multivariate fidelities [NMLW25], would presumably lead to different bounds.
Acknowledgements—ZH is supported by an ARC DECRA Fellowship (DE230100144) “Quantum-enabled super-resolution imaging” and an RMIT Vice Chancellor’s Senior Research Fellowship. She is also grateful to the Cornell School of Electrical and Computer Engineering for hospitality during an October 2025 research visit, as well as the Cornell Lab of Ornithology for the Merlin app, which has enabled many hours of fruitful birding. JJM acknowledges support from QuantERA (HQCC). TN acknowledges support from the Department of Mathematics and the IQUIST Postdoctoral Fellowship from the Illinois Quantum Information Science and Technology Center at the University of Illinois Urbana-Champaign. MMW acknowledges support from the National Science Foundation under grant no. 2329662 and from the Cornell School of Electrical and Computer Engineering.